### v1

# Nonlinear Permuted Granger Causality

Abstract
Granger causal inference is a contentious but widespread method used in fields ranging from economics to neuroscience. The original definition addresses the notion of causality in time series by establishing functional dependence conditional on a specified model. Adaptation of Granger causality to nonlinear data remains challenging, and many methods apply in-sample tests that do not incorporate out-of-sample predictability leading to concerns of model overfitting. To allow for out-of-sample comparison, we explicitly define a measure of functional connectivity using permutations of the covariate set. Artificial neural networks serve as featurizers of the data to approximate any arbitrary, nonlinear relationship, and under certain conditions on the featurization process and the model residuals, we prove consistent estimation of the variance for each permutation. Performance of the permutation method is compared to penalized objective, naive replacement, and omission techniques via simulation, and we investigate its application to neuronal responses of acoustic stimuli in the auditory cortex of anesthetized rats. We contend that targeted use of the Granger causal framework, when prior knowledge of the causal mechanisms in a dataset are limited, can help to reveal potential predictive relationships between sets of variables that warrant further study.
Keywords: Granger causality, permutation testing, out-of-sample predictability, multivariate time series, artificial neural networks
1 Introduction
Granger causal inference investigates the ability of a set of variables $\mathbf{X}\in\mathbb{R}^{T\times p}$ to predict future values of a response $\mathbf{Y}\in\mathbb{R}^{T\times d}$ in a multivariate time series, $t=1,\ldots,T$ (Granger, 1969; Wiener, 1956). The effect is traditionally measured through the variance of residuals in restricted and unrestricted models, like shown in Definition 1, where $\mathcal{P}(\mathbf{y}_{t}|\mathcal{I}_{<t})$ represents the optimal prediction of vector $\mathbf{y}_{t}$ given all information prior to time $t$. Subscripts indicating time constraints of the set of variables $\mathbf{X}$ are not included as we trivially assume that the covariates assigned to a vector response $\mathbf{y}_{t}$ precede time $t$.
Definition 1.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ if $\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\backslash\mathbf{X}\right)\right]$.
1.1 Discussion of Granger Causality
The form of Granger causality presented in Definition 1 is inherently conditional on additional information included in the modeling process. The optimal prediction $\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)$ is unattainable in practice, and the notion of Granger causality is a conditional model on some included explanatory covariate set $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the history of the response $\mathbf{Y}_{<t}$ like that in Definition 2.
Definition 2.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is conditionally Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the relevant history of the response $\mathbf{Y}_{<t}$ if $\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}\right)\right]$.
Model selection is implicit in the test for presence of Granger causality, and the inferential conclusion is coupled with a written form (Friston et al., 2013). Inclusion of additional variables strengthens the condition for establishing Granger causality; rejection of the null implies the covariate set $\mathbf{X}$ is found to provide unique and useful information for prediction of $\mathbf{y}_{t}$ beyond that contained in both $\mathbf{Z}$ and the lagged response (Granger, 1980). The basis for Granger causality requires fulfillment of several conditions including a sufficient length of continuous-valued stationary data, exact and complete specification of the model, error-free observation of the variables, and a sampling frequency on a regular discrete grid that contains the known lag relationship (Granger, 1969, 1980, 1988). In the form of Definition 2, we can relax the second condition provided that inference is accordingly narrowed to the conditional statement.
The framework can also be defined in terms of non-causality as a statement of conditional independence, where inclusion of additional variables $\mathbf{Z}$ is a simple extension (Florens and Mouchart, 1982; Granger, 1980).
Definition 3.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ does not Granger cause multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ if and only if $\mathbf{Y}_{<t+1}\perp\mathbf{X}$ given $\mathbf{Y}_{<t}$.
This statement is mathematically equivalent, but perhaps more powerful because it is defined in terms of the distributions of the variables, allowing for extension to several other forms of statistical tests. This definition may be prone to misuse if interpreted to place the burden of proof on establishing independence. In a definition from Section 4 of Shojaie and Fox (2022), time series $\mathbf{x}_{j}$ is Granger non-causal for time series $\mathbf{x}_{i}$ if and only if $\forall t$, $\mathcal{P}\left(\mathbf{x}_{ti}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<tj},\ldots,\mathbf{x}_{<tp}\right)=\mathcal{P}\left(\mathbf{x}_{ti}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<t(j-1)},\mathbf{x}_{<t(j+1)},\ldots,\mathbf{x}_{<tp}\right)$, or the equality of these predictions holds for all time points $t$. If misinterpreted to mean finding correlation at one time point in series $\mathbf{x}_{i}$ is enough to claim functional dependence of two time series, the presence of a causal connection effectively becomes the null hypothesis and Granger causality is reduced to an exceptionally weak statement.
Even when the question is framed with the onus on demonstrating a functional relationship, all Granger causal methods overreach their scope of reliable application when inference is performed on the individual variables included in the covariate set rather than on their collective behavior. Applications of the framework to interpreting individual model coefficients introduce a hidden multiplicity problem of repeated testing on subsets of $\mathbf{X}$, and the conclusion requires amendment to conditional non-causality of $\mathbf{x}_{i}$ given an exhaustive list of all other components $\mathbf{x}_{j}$, $j=1,,\ldots,i-1,i+1,\ldots,p$ in the model after adequate Type 1 error control. Methods that select causal covariate pairs through the use of penalized optimization problems don’t always allow for easy extension to the multiple testing problem, and may require repetitive sub-sampling approaches such as stability selection (Meinshausen and Bühlmann, 2010). A thorough definition of Granger causality provides a clear representation of the collective conclusion to be drawn on the covariate set $\mathbf{X}$, clarifies the conditional nature of the result on the specified model and included variables, and stresses the philosophical ordering from the null hypothesis implying no causal structure to the alternative that demands evidence of the contrary, all while retaining any general functional form of $\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\right)$.
We note the predictive nature of these definitions can relate to a causal structure between two variable groups, but alone is not enough to establish effective connectivity, distinguishing a direct influence of one population on another (Bressler and Seth, 2011; Friston, 1994). Even in the presence of the optimal set $\mathcal{I}_{<t}$, association and precedence are not enough to distinguish true causality if slight redundancies are included or an effect does not remain constant in direction through time (Maziarz, 2015). Granger causality exists in the realm of functional connectivity that identifies correlation at one or more time lags (Friston, 1994). Appropriate use of Granger causality is contentious, but the method has been applied to a variety of fields like economics, environmental sciences, and neuroscience (Bernanke, 1990; Cox Jr and Popken, 2015; Holland, 1986; Reid et al., 2019; Seth et al., 2015; Sims, 1972). Cautious and targeted use of the Granger causal framework can elucidate predictive relationships between variables that warrant further study when prior knowledge of potential causal relationships is limited.
In the linear realm, $\mathcal{P}(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X})$ is often sought from a vector autoregression (VAR) model, where models are evaluated with in-sample tests (Granger, 1969, 1980). Geweke (1982) proposed a spectral decomposition form of linear Granger causality for application to stationary Gaussian processes. Inference is performed with the estimated covariance of the restricted model (with $\mathbf{X}$ excluded) and that of the unrestricted model, where under the null it is assumed the two are equal. Increasing dimension of the response variable often requires implementation of an approximate test or a switch to permutation-like testing (Anderson and Robinson, 2001; Barnett and Seth, 2011). In-sample testing differs from the true notion of predictive ability, and out-of-sample methods align closer to the essence of Granger causality (Chao et al., 2001; Inoue and Kilian, 2005; Peters et al., 2016). This distinction is especially important when using deep learning techniques to model complex, nonlinear dynamics.
1.2 Nonlinear Adaptations
Nonparametric methods provide the basis for many nonlinear adaptations of Granger causality; specification of the exact functional form can be challenging. Parametric attempts, like the ordinary differential equations approaches of Henderson and Michailidis (2014) and Wu et al. (2014), allow for flexible definition of a series of functions to capture dependence, but are limited to modelling additive dynamics when the true mechanism of interaction may be more complicated. Some model-free information theoretic methods detect more elaborate forms of nonlinear dependence with minimal assumptions, but suffer from highly variable estimates and challenges of application to multivariate systems (Amblard and Michel, 2011; Runge et al., 2012; Vicente et al., 2011). Kernel Granger causality examines the linear form in a transformed feature space, but model comparison can become difficult (Marinazzo et al., 2008, 2011). Artificial neural networks allow for general forms of nonlinear dependence in a similar feature space.
Fully connected feed-forward networks (with $g$ a sigmoid function), like the construction shown in Equations 1 through 3, fall under the umbrella of the universal approximation theorems of Cybenko (1989) and Hornik (1991). These artificial networks act as featurizers of the data to a higher dimensional space, $\Psi:\mathbf{X}\in\mathbb{R}^{T\times p}\rightarrow\mathbf{H}\in\mathbb{R}^{T\times N}$, $N\gg p$; the architecture allows for arbitrarily close approximation $\hat{\mathbf{y}}_{t}$ of any function output $\mathbf{y}_{t}=f(\mathbf{x}_{t})$, provided $\mathbf{h}_{t}\in\mathbb{R}^{N}$ is allowed to contain a sufficient width of hidden states $N$ (Barron, 1993).
$\mathbf{h}^{1}_{t}$ $=g\left(\mathbf{W}^{1}\mathbf{x}_{t}+\mathbf{b}^{1}\right)$ (1)
$\mathbf{h}^{\ell}_{t}$ $=g\left(\mathbf{W}^{\ell}\mathbf{h}^{\ell}_{t}+\mathbf{b}^{\ell}\right)$ (2)
$\hat{\mathbf{y}}_{t}$ $=\mathbf{W}^{L}\mathbf{h}_{t}+\mathbf{b}^{L}$ (3)
In Equations 1 and 2, $g$ is an element-wise activation function, and $\ell=1,\ldots,L$ is the number of layers, or the “depth” of the network, where $\mathbf{W}^{1}\in\mathbb{R}^{N\times p}$, $\mathbf{W}^{\ell}\in\mathbb{R}^{N\times N}$, $\mathbf{W}^{L}\in\mathbb{R}^{d\times N}$, $\mathbf{b}^{\ell}\in\mathbb{R}^{N}$, and $\mathbf{b}^{L}\in\mathbb{R}^{d}$ are parameter matrices. The artificial networks shown above, sometimes called multilayer perceptrons (MLPs), and recurrent neural networks (RNNs) like long-short term memory networks (LSTMs) (Graves, 2012) and echo state networks (ESNs) (Jaeger, 2007) have all been used to investigate functional connectivity in time series data (Shojaie and Fox, 2022). RNN-type network construction adds a recurrence term to the hidden state form in Equation 1, $\mathbf{h}_{t}=g\left(\mathbf{W}^{r}\mathbf{h}_{t-1}+\mathbf{W}^{1}\mathbf{x}_{t}+\mathbf{b}\right)$.
Tank et al. (2022) extend Granger causality to the nonlinear space using component-wise MLP (cMLP) models, which model individual variables in the response $\mathbf{y}_{i}$, $i=1,\ldots,d$, with separate artificial networks. Parameters are sought via a penalized optimization approach and proximal gradient descent, and no Granger causal connection is inferred for an individual covariate if the corresponding row in a component input parameter matrix $\mathbf{W}^{1}_{i}$ is zero (Tank et al., 2022). The penalized approach encourages sparse solutions that block the inclusion of information from less predictive components in the hidden states $\mathbf{h}_{t}$, but selecting the regularization parameter is not an easy task and values may produce vastly different results. Tank et al. (2022) further introduce a component-wise LSTM (cLSTM) model that harnesses the recurrent structure to circumvent selection of the optimal lag for inclusion in the covariate set $\mathbf{X}$. This formulation, while making model specification as it relates to the time lag components easier, has the consequence of muddling inferential results across several lags. Khanna and Tan (2019) build on the framework of cMLP with statistical recurrent units, Biswas and Ombao (2022) discuss the application of the component network structure to frequency-specific relationships and non-stationary data, and Marcinkevics and Vogt (2021) introduce generalized vector autoregressive (GVAR) methodology aimed at interpretability of potential functional relationships. Other methods like the Jacobian Granger causality (JGC) method of Suryadi et al. (2023) use the Jacobian matrix, and Nauta et al. (2019) use convolutional neural networks (TCDF) and attention scores to serve as measures of variable importance.
Jointly estimating a large number of parameters is computationally expensive, and Duggento et al. (2021) instead use randomly initialized ESNs. The matrices $\mathbf{W}^{1}$ and $\mathbf{b}^{1}$ are populated with random values, and learning only takes place in the output weights $\mathbf{W}^{L}$ via penalized regression (Jaeger, 2007). Because computation is performed using linear techniques rather than a gradient descent algorithm, complexity decreases; however, inference can only be performed on the output coefficients if information mixing does not occur in the hidden states $\mathbf{h}_{t}$. They formulate $\mathbf{W}^{1}$ as a block diagonal matrix, which limits the scope of application to a specific subset of additive nonlinear interactions (Duggento et al., 2021).
Many of these methods ignore the collective inference principle of Granger causality and instead take the eager approach of performing individual covariate inference, sometimes with disregard for the multiplicity problem. Evaluation of these predictive relationships is often performed using in-sample tests. With their universal approximator ability, artificial neural networks of sufficient width or depth can approximate any functional relationship between two covariate sets, even if it is data-specific and the model is overfit, making them prone to link variables that do not have a predictive relationship as the dimension of the network increases. Sparsity inducing penalties may marginally improve reliability of in-sample tests, but out-of-sample testing helps control the overfitting problem to identify only useful functional relationships.
In this vain, Horvath et al. (2022) develop the Learned Kernel VAR (LeKVAR) method that proposes use of a kernel parameterized by an artificial neural network they argue is less prone to overfitting from a decoupling importance measure of the individual series and the selected lags, and the TCDF method employs a permutation-like testing procedure after variable selection (Nauta et al., 2019). Our goal is to formulate a simple method for identification of nonlinear functional relationships that lends itself to out-of-sample testing and clear control of the multiplicity problem in a Granger causal framework.
1.3 Overview
Rather than comparison of the inherently unequal restricted and unrestricted model errors, we focus out-of-sample predictability by implementing a permutation structure. There is precedence for the use of permutation-type procedures on general linear models, and their asymptotics are well studied in literature (Anderson and Legendre, 1999; Anderson and Robinson, 2001; DiCiccio and Romano, 2017; Winkler et al., 2014). Nauta et al. (2019) employ this type of importance measure after their complicated CNN screening procedure, and we argue for its widespread use as a decision framework in Granger causal inference. Importance of a chronologically ordered variable can be interpreted as “causal” (predictive) effect, and the strategy builds on the concept of exchangeability (like the directed graph method of Caron and Fox (2017)), where if $\tilde{\mathbf{X}}$ is a random permutation of the rows of $\mathbf{X}$, $\mathbf{Y}_{<t+1}\perp\mathbf{X}$ given $\mathbf{Y}_{<t}$ implies $\mathbf{Y}_{<t+1}\perp\tilde{\mathbf{X}}$ given $\mathbf{Y}_{<t}$, but the converse is not always true (Van der Laan, 2006).
To investigate if $\mathbf{X}$ Granger causes $\mathbf{Y}$, we propose the following definition pair, adjusting Definitions 2 and 3 to a permutation structure for the covariate matrix. The unrestricted and restricted models are replaced by a null model and a permuted model.
Definition 4a.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is not conditionally Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the relevant history of the response $\mathbf{Y}_{<t}$ if and only if $\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\overset{d}{=}\mathbf{Y}_{<t+1}|\mathbf{Y}_{<t},\mathbf{Z},\tilde{\mathbf{X}}$, where $\tilde{\mathbf{X}}$ is a copy of $\mathbf{X}$ with the time axis (rows) randomly permuted.
Definition 4b.
[…]
The permutations of $\mathbf{X}$ break the potential dependence structure with the response while retaining the intradependence of the covariates, as information in a given row across columns remains intact. The restructuring of the Granger causal framework allows for use of out-of-sample estimated prediction errors, corrects the imbalance of comparison between restricted and unrestricted, and presents a clear path to account for multiple testing, aligning the methodology closer to its inferential utility.
In Section 2, we outline the nonlinear permuted Granger causal methodology (NPGC) as an alternative framework to penalized optimization approaches, and we discuss the theoretical results in Section 3. Section 4 presents performance for the permutation method for simulated data, Section 5 applies NPGC to publicly available whole-cell recordings from the primary auditory cortex of an anesthetized rat in response to natural sound fragments, and Section 6 concludes the manuscript.
2 Methodology
Suppose observed realizations of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ arise from the set of all potential realizations $\omega\in\Omega$. We define $\Omega_{\text{obs}}$ as the size $\varphi$ set of observations, $\Omega_{\text{obs}}=\left\{1,\ldots,\varphi\right\}\subseteq\Omega$, and note that usually $\varphi=1$. Instances for $\varphi>1$ may occur with repeated trials of a controlled experiment. For simplicity of the original presentation, we omit the subscript $\omega$ notation specifying an observed realization until the end of this subsection.
With familiarity of a dataset, a researcher selects $\gamma$ lagged values of $\mathbf{Y}$ that serve as a representative history, $\mathbf{Y}_{\text{lag}}\in\mathbb{R}^{(T+\gamma)\times\gamma d}$ (corresponding to $\mathbf{Y}_{<t}$ in Definitions 4a and 4b), and assume that an appropriate number of lags is selected such that autocorrelation in $\mathbf{Y}$ is fully explained across all potential realizations. Selection of the truncation lag $\gamma$ is outside the scope of this work, and we refer the reader to Ng and Perron (2001), Ivanov and Kilian (2005), Shojaie and Michailidis (2010), and Nicholson et al. (2017) for detailed discussions. With a finite data length $T+\gamma$, this process restricts the usable portion of $\mathbf{X}$, $\mathbf{Y}$, and any additional covariates $\mathbf{Z}$ to the last $T$ rows. All columns are standardized for consistent behavior in the featurization process with an activation function (Goodfellow et al., 2016).
The rows of the covariate matrix $\mathbf{X}$ are randomly reorganized via $\bm{\Pi}_{m}$ to generate several permutations $\tilde{\mathbf{X}}_{m}=\bm{\Pi}_{m}\mathbf{X}$ for $m=1,\ldots,M$. The designated first permutation, $m=1$, corresponds to the original ordering of the data where $\bm{\Pi}_{1}=\mathbf{I}$. Dependence across rows of $\mathbf{X}$, for example a covariate lag structure, is captured by augmenting the matrix with additional columns.
After permutation, we compile the predictor matrices $\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\in\mathbb{R}^{T\times(1+\gamma d+q+p)}$, and set the nonlinear activation function $g=\tanh$ in the functional transformation $\Psi$ of Equation 1. With the inclusion of the intercept term, $\Psi$ becomes
$\mathbf{H}_{m}$ $=\Psi\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\right)=\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\tilde{\mathbf{X}}_{m}\right]\mathbf{W}\right),$
$\text{where}\quad\mathbf{W}$ $=(w_{ij})\in\mathbb{R}^{(1+\gamma d+q+p)\times N},\quad w_{ij}\sim\mathcal{N}(0,1).$ (4)
The transformation matrix is generated via random initialization as indicated. We account for added uncertainty in these random models by generating several initializations $\mathbf{W}_{r}$ for $r=1,\ldots,\mathscr{R}$ and extracting the aggregate behavior. We can write the models in terms of the original ($m=1$) and permuted ($m=2,\ldots,M$) feature spaces: $\mathbf{Y}=\mathbf{H}_{m,r}\mathbf{B}_{m,r}+\mathbf{U}_{m,r}$ (where $\mathbf{B}_{m,r}$ corresponds to $\mathbf{W}^{L}$ in Equation 3). The exact formulation of $\Psi$ in Equation 4 is not the main focus of this manuscript and the dimension of the feature space $N$ is fixed for direct comparison to other methods. For demonstration of our method, we select a simple, widely-applicable model agnostic structure, and discuss potential adaptations in the supplement, including a cross validation approach to select an optimal dimension.
Define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the prediction for $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given the relevant history of the response $\mathbf{Y}_{\text{lag}}$, the additional variables $\mathbf{Z}$, and the permuted covariate set $\tilde{\mathbf{X}}_{m}$. We denote $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$ as the corresponding parameter over all potential realizations $\omega\in\Omega$. Variation in the estimate arises from potential realizations of the data $\omega\in\Omega$ and via random initializations approximating the nonlinear functional form. For a given data realization $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, and under the true functional form, we define the specific covariance matrix of the prediction $\bm{\Sigma}_{m,\omega}$; for random initialization $r$, we define the covariance matrix of prediction $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$. We estimate the out-of-sample variation parameter $\vartheta_{m}$ for each permutation with a cross validation approach.
2.1 Estimating Granger Causal Influence
A sufficiently large value featurization dimension $N$ is chosen to “linearize” any existing functional relationship, but not so large that the network is able to memorize inputs or fabricate dependence between the permuted data and a response. We implicitly assume the existence of some nonlinear functional relationship between a covariate set and a response will be “easier” for an artificial network to learn than random matching of inputs to outputs in the permuted data, and propose an exact form of this condition in Section 3. The data is split into $K$ sets for model computation and testing, and we define the number of observations in each set $k=1,\ldots,K$ as $T_{k}=\lfloor T/K\rfloor+\mathbf{1}\left\{\left(T\bmod K\right)\geq k\right\}$.
We generate $r=1,\ldots,\mathscr{R}$ featurizations under the form of Equation 4. The model matrices $\mathbf{W}_{r}$ are held fixed over all permutations $m$ and observations $\omega\in\Omega_{\text{obs}}$ for a consistent error estimation framework. For each permutation, with $m=1$ corresponding to the original data, the predictor matrices are projected into the respective feature spaces, and we can write the residuals for test set $k$ as
$\mathbf{R}_{\omega,m,r,k}$ $=\mathbf{H}_{\omega,m,r,k}\left(\mathbf{H}_{\omega,m,r,-k}^{\prime}\mathbf{H}_{\omega,m,r,-k}\right)^{-1}\mathbf{H}_{\omega,m,r,-k}^{\prime}\mathbf{Y}_{\omega,-k},-\mathbf{Y}_{\omega,k}$ (5)
where the training data (subscript k-k) excludes set $k$.
These out-of-sample prediction residuals from the test set align closer to the original definition of predictive ability in Granger causal inference than the in-sample model variation. This distinction is especially important with the use of artificial networks and the ability to learn any arbitrary, data-specific dependence structure. We obtain the estimate for the out-of-sample variation in prediction residuals $\hat{\vartheta}_{m}$ with Equation 6.
$\hat{\vartheta}_{m}$ $=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}_{\omega,m,r,k}^{\prime}\mathbf{R}_{\omega,m,r,k}\right)$ (6)
Using the permutation structure, we approximate a null distribution of variation from each model $m=1,\ldots,M$. These random permutations of the covariate set $\mathbf{X}$ break the potential dependence structure present in the form of some predictive relationship with the response $\mathbf{Y}$. Under the null hypothesis, we can view the original data as one of $M$ random permutations. We expect that the original “permutation” $\tilde{\mathbf{X}}_{1}=\mathbf{X}$ will exhibit similar properties to $\tilde{\mathbf{X}}_{m}$ for $m=2,\ldots,M$, or if the time observations $\mathbf{X}$ are exchangeable for prediction of $\mathbf{Y}$, the conditional distribution of the prediction does not change. The variation estimates are drawn from a distribution of all possible permutations,
$\hat{\vartheta}_{m}\sim\hat{\mathcal{H}}(s)$ $=(T!)^{-1}\sum_{i=1}^{T!}\mathbf{1}\{\hat{\vartheta}_{i}\leq s\},$ (7)
and we expect $\hat{\vartheta}_{1}$ to fall above some lower tail portion. This leads to an approximate null distribution where the estimate $\hat{\vartheta}_{1}$ is at quantile $\hat{Q}_{M}$ of the empirical distribution $\hat{\mathcal{H}}_{M}(s)$, defined in Equation 8, formed from a subsample of size $M\leq T!$. A decision rule is formulated from comparison to a chosen level of test $\alpha$.
$\hat{\mathcal{H}}_{M}(s)$ $=\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\{\hat{\vartheta}_{m}\leq s\}$ (8)
$\hat{Q}_{M}=\hat{\mathcal{H}}_{M}(\hat{\vartheta}_{1})$ $=\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\hat{\vartheta}_{m}\leq\hat{\vartheta}_{1}\right\}$ (9)
Rejection of the null hypothesis in this framework, $\hat{Q}_{M}\leq\alpha$, presents evidence for $\mathbf{X}$ as Granger causal of $\mathbf{Y}$ conditional on the additional variables $\mathbf{Z}$ and the relevant history of the response $\mathbf{Y}_{\text{lag}}$. For the case when $\varphi=1$, we also clarify that this conclusion is conditional on error free observation of the dataset. Inferential results must either include this assumption, or the scope narrowed to the observation $\omega$.
The full process is given in the form of pseudocode in the supplementary material. We explicitly outline the theoretical behavior of these estimates and the development of the underlying framework in Section 3.
3 Theory
We define $\bm{\Theta}_{m}$ as the underlying covariance matrix of the predictive ability of permutation $m$, and the quantity $\vartheta_{m}=\text{tr}\left(\bm{\Theta}_{m}\right)$. For each potential realization of the data $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$, $\omega\in\Omega$, we define the realization-specific covariance matrix $\bm{\Sigma}_{m,\omega}$ that is drawn from some distribution with expectation $\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]=\vartheta_{m}$ and variance $\tau_{\omega}^{2}<\infty$ that is constant over all permutations. Each random initialization for the featurization process $r=1,\ldots,\mathscr{R}$ produces $\mathbf{S}_{m,\omega,r}$ as an estimate of $\bm{\Sigma}_{m,\omega}$, where $\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and $\text{Var}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\tau_{r}^{2}<\infty$.
Using the permutation structure, we evaluate the null and permuted models with the out-of-sample prediction residuals, shown in Equation 5, and define the estimate for total variation in Equation 6. Under the null hypothesis when the conditional distribution of the prediction is invariant to permutation of the covariate set $\mathbf{X}$, $\vartheta_{1}=\vartheta_{2}=\cdots=\vartheta_{T!}=\vartheta$, and we write the null and alternative hypotheses below.
$H_{0}$ $:\vartheta_{1}=\vartheta_{2}=\cdots=\vartheta_{T!-1}=\vartheta_{T!}$
$H_{A}$ $:\vartheta_{1}<\vartheta_{i}\quad\text{for all $i=2,\ldots,T!$}$ (10)
We test the null hypothesis in Equation 10 using the sample quantile $\hat{Q}_{M}$ from the empirical distribution $\hat{\mathcal{H}}_{M}(s)$ defined in Equations 8 and 9, and outline theoretical results under a set of mild conditions that are comparable to those found in relevant literature. Four additional conditions in the supplement provide regularity to the featurization process.
Condition 1.
The data is continuous and stationary, and the discrete, regular sampling grid $t=1,\ldots,T$ is sufficiently fine to capture any potential functional dependence in the variable matrices $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for any realization $\omega\in\Omega$.
Condition 2.
For all realizations $\omega\in\Omega$, and random initializations $\mathbf{W}_{r}\in\mathcal{W}$, $r=1,\ldots,\mathscr{R}$, the quantities $\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}$ are drawn i.i.d. from continuous distributions with $\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]=\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and constant variance across permutations and potential realizations $\text{Var}\left(\text{tr}\left[\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)=\tau_{r}^{2}<\infty$. Similarly, $\bm{\Sigma}_{m,\omega}$ are i.i.d. from continuous distributions with expectation $\mathbb{E}\left[\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]=\vartheta_{m}$ and constant variance across permutations $\text{Var}\left(\text{tr}\left[\bm{\Sigma}_{m,\omega}\right]\right)=\tau_{\omega}^{2}<\infty$.
Condition 3.
The relevant history of the response, $\mathbf{Y}_{<t,\omega}=\mathbf{Y}_{\text{lag},\omega}\in\mathbb{R}^{T\times\gamma d}$ is appropriately chosen such that the model errors $\mathbf{U}_{m,\omega,r}$ are independent, or $\mathbf{u}_{m,\omega,r,t}\perp\mathbf{u}_{m,\omega,r,t^{\prime}}$ for any realization $\omega\in\Omega$, permutation $m=1,\ldots,M$, initialization $r=1,\ldots,\mathscr{R}$, and time point $t=1,\ldots,T$ where $t^{\prime}\neq t$. Further, the model errors follow multivariate normal distributions with mean zero and constant variation, leading to the result
$\mathbf{u}_{m,\omega,r,t}$ $\overset{i.i.d.}{\sim}\mathcal{N}_{d}\left(\mathbf{0},\mathbf{S}_{m,\omega,r}\right).$ (11)
3.1 Asymptotic Properties
We examine the asymptotic behavior of the estimates and established testing framework under the listed conditions in this section and the supplement. Define the underlying variation parameter for permutation $m$ as $\vartheta_{m}$, and the estimate $\hat{\vartheta}_{m}$ as in Equation 6. We assume a fixed test set size $T_{k}$ and allow the training set $T_{-k}=T-T_{k}$ and number of folds $K$ to grow as $T\rightarrow\infty$.
Theorem 1.
Under the conditions listed in Section 3 and the supplement, with Condition 2 modified such that $\tau_{\omega}^{2}=0$ (i.e., we achieve error free observation of the data), for all $\varepsilon>0$,
$\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(|\hat{\vartheta}_{m}-\vartheta_{m}|\leq\varepsilon\right)=1.$ (12)
Alternatively, under the conditions listed in Section 3 and the supplement, for all $\varepsilon>0$,
$\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(|\hat{\vartheta}_{m}-\vartheta_{m}|\leq\varepsilon\right)=1.$ (13)
Without error free observation of the data, the underlying variation $\tau_{\omega}^{2}$ remains present in the estimate, but shrinks as the number of observations gets large.
Theorem 2.
Under the conditions listed in Section 3 and the supplement, the estimate for the variation parameter $\hat{\vartheta}_{m}$ admits a Central Limit Theorem with respect to the number of observations $\varphi$ in $\Omega_{\text{obs}}$.
$\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\sqrt{\varphi}\left(\hat{\vartheta}_{m}-\vartheta_{m}\right)$ $\xrightarrow{d}\mathcal{N}\left(0,\tau_{\omega}^{2}\right)$ (14)
As a direct result of Theorem 2, $\hat{\vartheta}_{m}$ is a consistent estimate for $\vartheta_{m}$ when $\tau_{\omega}^{2}>0$ as $\varphi$ tends to infinity along with the size of the training and test sets. Next, we examine the test statistic $\hat{Q}_{M}$ under the null and alternative hypotheses.
3.1.1 Under the Null Hypothesis
Under the null hypothesis, the underlying variation parameter is constant for every possible permutation, $\vartheta_{1}=\cdots=\vartheta_{T}!$. We define the quantile estimate $\hat{Q}_{M}$ as in Equation 9, and can establish the limiting uniform distribution from the result of Theorem 1.
Theorem 3.
Under the null hypothesis and the conditions listed in Section 3 and the supplement,
$\lim_{T\rightarrow\infty}\lim_{M\rightarrow T!}\hat{Q}_{M}\xrightarrow{d}\text{Uniform}(0,1).$ (15)
3.1.2 Under the Alternative Hypothesis
We define $Q$ as the true quantile for parameter $\vartheta_{1}$ over all $T!$ possible permutations. In the limit, the quantile estimate defined in Equation 9 converges in probability to $Q$.
Theorem 4.
Under the alternative hypothesis and the conditions listed in Section 3 and the supplement, with Condition 2 modified such that $\tau_{\omega}^{2}=0$ (i.e., we achieve error free observation of the data), for all $\varepsilon>0$,
$\lim_{M\rightarrow T!}\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(|\hat{Q}_{M}-Q|\leq\varepsilon\right)=1.$ (16)
Similarly, under the alternative hypothesis and the conditions listed in Section 3 and the supplement, for all $\varepsilon>0$,
$\lim_{M\rightarrow T!}\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(|\hat{Q}_{M}-Q|\leq\varepsilon\right)=1.$ (17)
4 Simulation Study
We evaluate the ability of the permutation-based methodology of NPGC (similar to the decision rule formulation of TCDF) to detect the presence of functional connectivity and control for false positive results with a consistent single-layer cMLP nonlinear transformation framework. All methods use featurization dimension $N=100$. Three of the five comparison methods use in-sample methodology and variants of the lasso penalty within cMLP models (Tank et al., 2022; Tibshirani, 1996). The other two methods employ random initialization like NPGC, and use in-sample testing and decision rules from comparison of restricted and unrestricted models or the substitution of random Gaussian noise.
Multiple testing on bivariate pairs is not examined; we consider a group decision process where collectively the covariate set $\mathbf{X}$ is (or is not) Granger causal for the response $\mathbf{Y}$. Ability to detect a Granger causal result will almost certainly improve with a more sophisticated nonlinear structure; we demonstrate the relative differences between methods at a baseline level. A full list of included methods is given in Table 1.
(i) Nonlinear Permuted Granger Causaltiy (NPGC)
(ii) cMLP, Group Lasso Penalty (cMLP-GL)
(iii) cMLP, Group Sparse Group Lasso Penalty (cMLP-GSGL)
(iv) cMLP, Hierarchical Lasso Penalty (cMLP-H)
(v) Restricted vs. Unrestricted Models (R/U)
(vi) Gaussian Noise Substitution (GNS)
Table 1: Testing frameworks for Granger causal inference simulations.
The lasso type objectives of cMLP methods (ii) - (iv) penalize nonzero rows of the matrix $\mathbf{W}^{1}$ in the formulation of Equation 1. The corresponding null and alternative hypotheses are adjusted to those shown in Equation 18, and a Granger causal connection is present for a covariate set if any corresponding rows of $\mathbf{W}^{1}$ contain nonzero components (Tank et al., 2022).
[…]
A group lasso penalty is applied to each variable $i$ in the model matrix $\mathbf{W}^{1}$ corresponding to the covariate set $\mathbf{X}$ over all time lags $t,\ldots,t-\gamma$ (Yuan and Lin, 2006). As in Tank et al. (2022), we define $\mathbf{W}^{1}_{i}=\left[\mathbf{W}^{1}_{i,t}\;\cdots\;\mathbf{W}^{1}_{i,t-\gamma}\right]$, with each entry as the weights of the model matrix row for variable $i$ and lagged data up to $t-\gamma$. Penalties for the cMLP methods take the general form $\lambda\sum_{i=1}^{p}\beta\left(\mathbf{W}^{1}_{i}\right)$, with specific $\beta$ for each defined below.
$\beta_{GL}\left(\mathbf{W}^{1}_{i}\right)$ $=\left\|\mathbf{W}^{1}_{i}\right\|_{F}$
$\beta_{GSGL}\left(\mathbf{W}^{1}_{i}\right)$ $=\left\|\mathbf{W}^{1}_{i}\right\|_{F}+\sum_{k=0}^{\gamma}\left\|\mathbf{W}^{1}_{i,t-k}\right\|_{2}$
$\beta_{H}\left(\mathbf{W}^{1}_{i}\right)$ $=\sum_{k=0}^{\gamma}\left\|\left[\mathbf{W}^{1}_{i,t-k}\;\cdots\;\mathbf{W}^{1}_{i,t-\gamma}\right]\right\|_{F}$ (19)
The group sparse group lasso penalty combines sparsity of included variables and their lagged values like in Simon et al. (2013), and the novel hierarchical penalty of Tank et al. (2022) retains information about the natural ordering of the variables, encouraging solutions where for some lag $^{*}, k>k$$k^{*}$ implies $\mathbf{W}^{1}_{i,t-k}=\mathbf{0}$. We select a regularization parameter of $\lambda=0.5$ based on a cross-validation like trial and error approach. Conclusions drawn from these methods can vary greatly depending on the chosen $\lambda$; smaller values retain all matrix entries and larger values penalize the matrix to zero.
We formulate two naive methods for comparison to NPGC. The in-sample restricted and unrestricted comparison examines the ratio of model residuals $\hat{\sigma}^{2}_{\text{Res}}/\hat{\sigma}^{2}_{\text{Unres}}$ in a randomly initialized cMLP model. The in-sample Gaussian noise substitution is methodologically similar, but instead substitutes Gaussian noise in place of the covariate set $\mathbf{X}$ to examine $\hat{\sigma}^{2}_{\text{Noise}}/\hat{\sigma}^{2}_{\text{Unres}}$. The three model errors are shown in Equation 20, where $\mathbf{E}\sim\mathcal{MN}_{T\times p}\left(\mathbf{0},\mathbf{I},\mathbf{I}\right)$, and the corresponding $\hat{\sigma}^{2}$ shown in Equation 21.
$\mathbf{U}_{\text{Unres},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{X}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Res},r}$
$\mathbf{U}_{\text{Res},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{0}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Res},r}$
$\mathbf{U}_{\text{Noise},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{E}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Noise},r}$ (20)
$\hat{\sigma}_{r}^{2}$ $=T^{-1}\mathbf{U}_{r}^{\prime}\mathbf{U}_{r}$ (21)
The individual terms in Equation 21 follow chi-square distributions with degrees of freedom TNT-N, and each ratio for an individual initialization is distributed $\mathcal{F}_{T-N,T-N}$. Over all initializations $r=1,\ldots,\mathscr{R}$ the sum of a large number of these independent (via Condition 3) statistics is approximately normal. As $\mathscr{R}\rightarrow\infty$ under the null hypothesis,
$\mathscr{R}^{-1}\sum_{r=1}^{\mathscr{R}}\frac{\hat{\sigma}_{0,r}^{2}}{\hat{\sigma}_{1,r}^{2}}\xrightarrow{d}\mathcal{N}\left[\frac{T-N}{T-N-2},\frac{4(T-N)^{2}(T-N-1)}{\mathscr{R}(T-N)(T-N-2)^{2}(T-N-4)}\right].$ (22)
We formulate a naive decision rule with this limiting distribution and a large number ($\mathscr{R}=1000$) of random initializations.
The NPGC methodology is performed with $M=400$ permutations, $K=5$ cross validation folds, and $\mathscr{R}=50$ random initializations. We evaluate the methods based on correctly flagging a Granger causal result when direct functional dependence is present, or correctly labelling a non-causal result when it is absent, leading to the potential outcomes in Table 2. For the NPGC and naive ratio methods, we flag a result if the quantile estimate of the associated statistic under the null hypotheses is under a specified level $\alpha$. The lasso-type methods lack the direct translation to a traditional hypothesis testing framework.
Result $\backslash$ Truth Granger causal ($\text{GC}=1$) Not causal ($\text{GC}=0$)
Granger causal ($\text{GC}=1$) $\rho_{1}=\sum_{\mathcal{S}}\mathbf{1}\left\{\mathcal{D}=1\right\}/N_{1}$ $\rho_{10}=\sum_{\mathcal{S}}\mathbf{1}\left\{\mathcal{D}=1\right\}/N_{0}$
Not causal ($\text{GC}=0$) $\rho_{01}=\sum_{\mathcal{S}}\mathbf{1}\left\{\mathcal{D}=0\right\}/N_{1}$ $\rho_{0}=\sum_{\mathcal{S}}\mathbf{1}\left\{\mathcal{D}=0\right\}/N_{0}$
Table 2: Proportion of potential outcomes for the decision $\mathcal{D}$ of each of the methods shown in Table 1. The set $\mathcal{S}$ represents the space of all simulations (both $\text{GC}=1$ and $\text{GC}=0$) within a given setting, $N_{1}=\sum_{\mathcal{S}}\mathbf{1}\left\{\text{GC}=1\right\}$, and $N_{0}=\sum_{\mathcal{S}}\mathbf{1}\left\{\text{GC}=0\right\}$.
4.1 Simulation Settings
We test NPGC and the comparator methods in Table 1 on two nonlinear processes: Lorenz-96 models of Karimi and Paul (2010), and threshold autoregressive models (TAR) introduced by Tong (1978). The $p$-dimensional Lorenz-96 model ($p\geq 4$) is governed by the continuous differential equation
$\frac{dx_{i,t}}{dt}$ $=\left(x_{i+1,t}-x_{i-2,t}\right)x_{i-1,t}-x_{i,t}+F,$ (23)
with $i=1,\ldots,p$ and boundary series $x_{-1,t}=x_{p-1,t}$, $x_{0,t}=x_{p,t}$, and $x_{p+1,t}=x_{1,t}$. $F$ is a forcing constant we generate as $F\sim\mathcal{U}(5,20)$ with higher values introducing a larger degree of nonlinear, chaotic behavior. For data generation, we use a sampling rate of $\Delta t=0.05$ and allow a burn-in period of 500 time steps.
The TAR(2) model of Tong (1978) is governed by a similar skeleton to a vector autoregressive model, but allows for changes in parameters based on the value of a threshold variable (Tong and Lim, 1980).
$\begin{bmatrix}\mathbf{x}_{t}\\ \mathbf{x}_{t-1}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{A}^{(k)}_{1}&\mathbf{A}^{(k)}_{2}\\ \mathbf{I}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{x}_{t-1}\\ \mathbf{x}_{t-2}\end{bmatrix}+\begin{bmatrix}\bm{\varepsilon}^{(k)}_{t}\\ \mathbf{0}\end{bmatrix}$. (24)
We define the threshold on the values in $\mathbf{x}_{t-2}$ with regime $k=1$ corresponding to the case when $\sum_{i=1}^{p}x_{i,t-2}\leq 0$ and $k=2$ to $\sum_{i=1}^{p}x_{i,t-2}>0$. For the TAR(2) process, $\bm{\varepsilon}^{(k)}_{t}\sim\mathcal{N}_{p}\left(\mathbf{0},\sigma_{k}^{2}\mathbf{I}\right)$ where $\sigma_{1}=0.5$ and $\sigma_{2}=0.2$. Each $\mathbf{A}^{(k)}_{1},\mathbf{A}^{(k)}_{2}\in\mathbb{R}^{p\times p}$ has elements $\mathbf{A}^{(k)}=(a^{(k)}_{ij})\sim\mathcal{U}(-0.5,0.5)$ that are thresholded to zero if $|a^{(k)}_{ij}|<0.1$, and their spectral radii are at most 0.8 to ensure stationary data generation. From the initial zero state, we allow for a burn-in period of 500 time steps as in the Lorenz-96 data generation procedure.
We generate two groups of dependent data observations with $p=6$ (one from each process), containing $\mathbf{x}_{i}$ for $i=1,\ldots,12$. We label $i=1,\ldots,6$ as the Lorenz-96 process and $i=7,\ldots,12$ as the TAR(2) process. The samples are of length $T=250$, 500500, and $1000$ after burn-in and truncation for lag selection to include in the model, $\gamma=3$. We select one index to serve as the response, three to serve as the set of additional variables $\mathbf{Z}$ and three for the covariate set $\mathbf{X}$ depending on a Granger causal designation. A dataset is designated “Granger causal” if at least one series in $\mathbf{X}$ has a direct influence on the selected response (i.e., one series in $\mathbf{X}$ is chosen from the same generating process). We simulate 200 datasets for each of the settings shown in Table 3 for a total of 4800 trials.
Response Variable $T$ # Causal in $\mathbf{X}$ (GC) # Causal in $\mathbf{Z}$
Lorenz-96 $\left\{250,500,1000\right\}$ $\left\{0,2\right\}$ $\left\{0,2\right\}$
TAR(2) $\left\{250,500,1000\right\}$ $\left\{0,2\right\}$ $\left\{0,2\right\}$
Table 3: Settings for simulation and selection of variables. A variable is labelled causal if from the same generating process as the response $\mathbf{Y}$. Variables are randomly selected from within their causal or non-causal groups.
4.2 Simulation Results
We present simulation results split by Granger causal designation and process of the chosen response variable. For each, we examine the control for false positive results and the ability to detect a Granger causal connection. We do not examine AUROC as it does not implicate a decision rule a priori. Tables 4(a) and 4(b) list the proportion of correctly identified causal relationships $\rho_{1}$ and Figures 1 and 2 the observed Type 1 Error by specified level of test. The results for the finer grid of designations from Table 3 are given in the supplement.
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 0.915 0.870 0.155 0.555 0.752 0.989
$T=500$ 0.965 0.852 0.132 0.558 0.842 0.998
$T=1000$ 0.985 0.850 0.122 0.500 0.850 1.000
(a) TAR(2) response
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 0.978 0.765 0.232 0.482 0.025 1.000
$T=500$ 0.988 0.728 0.160 0.442 0.001 1.000
$T=1000$ 0.992 0.715 0.128 0.412 0.020 1.000
(b) Lorenz-96 response
Table 4: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2). Results are presented by the fine grid of designations in Table 3 in the supplementary material.
The NPGC method succeeds in identifying many cases of Granger causal influence in both processes, and as expected, this ability approaches the upper limit of one as the number of time points increases. The cMLP methods provide spotty results, and these are greatly influenced by the selected parameter $\lambda$. Without guidance for a specific penalty selection, the global decision of “Granger causal” or “non-causal” is uncertain. The naive restricted and unrestricted model method fails to identify many Granger causal pairings in the Lorenz-96 trials. The Gaussian noise substitution method correctly identifies nearly all Granger causal pairs, but this comes at the cost of uncontrolled Type 1 error (see Figures 1 and 2).
Figure 1: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations. Methods that adhere to a chosen level of test will lie closer to the uniform cdf (gray) included for reference.
Figure 2: Type 1 Error control by reference quantile $\alpha$ for Lorenz-96 simulations. Methods that adhere to a chosen level of test will lie closer to the uniform cdf (gray) included for reference.
Under the null hypothesis, we demonstrate moderate adherence to the asymptotic properties shown in Theorem 3 by presenting the observed Type 1 Error control by generic reference quantile $\alpha$. The NPGC permutation methodology tracks relatively close to the included uniform cdfs in Figure 1 for TAR(2) simulation settings, and exhibits mild to moderate deviations for Lorenz-96 settings in Figure 2. We can likely attribute these deviations to the complicated nature of the chaotic Lorenz system; we do not see this same pattern in other simulations when both groups are generated by a TAR(2) process (see Figure 7 in the supplement).
The cMLP lasso methods are included as horizontal lines in these plots; they do not have a straightforward extension for a specific level of test, other than computationally expensive iteration over several penalties $\lambda$. The naive methods do not adhere to their theoretical Type 1 error control, but approaches in the noise substitution realm appear promising if this issue can be resolved.
Granger causal identification methods that use a penalized objective can be useful, but do not allow for the global decision to label a group of variables “Granger causal” or “non-causal”. These methods should be applied to narrow the scope of research to individual variables of the covariate set after a global method is applied.
5 Application Study
The out-of-sample permutation framework is applied to neuronal responses to acoustic stimuli in the primary auditory cortex of an anesthetized (ketamine–medetomidine) rat. The data, taken from the Collaborative Research in Computational Neuroscience data sharing website, consists of in vivo whole-cell recordings (mV) sampled at 4kHz in response to natural sound fragments (Asari et al., 2009; Machens et al., 2004). We examine experimental trial 60 from the 050802mw03 data and isolate a time region of interest between 11 and 15 seconds (partially shown in Figures 2D and 2E of Machens et al. (2004)). The subset of data corresponds to the recordings in response to a jaguar (Panthera onca) mating call sound fragment presented at 97.656 kHz; we resample to a frequency of 4kHz, matching that of the response. To construct a simplistic non-causal scenario, we include additional sound fragments of a Humpback whale (Megaptera novaeangliae) and Knudsen’s frog (Leptodactylus knudseni) that are played in other trials throughout the experiment. Figure 3 displays the whole-cell recordings to the jaguar mating call and all sound fragments examined. We refer to Machens et al. (2004) for additional detail on data collection and experimental methods.
Figure 3: In vivo whole-cell recordings (mV) from the contralateral (left) primary auditory cortex of an anesthetized rat (top) in response to a jaguar mating call acoustic stimulus fragment resampled at 4kHz (bottom). Humpback whale and Knudsen’s frog acoustic stimuli are included as non-causal examples for the methodology.
We examine the performance of the Granger causal detection methodology by formulating a farcical example in which any suitable method should be able to recover the correct causal structure. The methods are expected to flag the jaguar sound fragment as causal, while correctly labelling the other two included fragments non-causal. We isolate lagged information from 30ms to 40ms, consistent with the measured the half-maximal synaptic conductance in Wehr and Zador (2003). One acoustic stimulus fragment is chosen as the covariate set $\mathbf{X}$ and the others are additional variables $\mathbf{Z}$ in the model. Lagged values of the response are included up to 10ms. We implement the NPGC method with $M=400$ permutations, $K=5$ cross validation folds, $\mathscr{R}=50$ random initializations, and due to the large number of variables, we increase the feature dimension to $N=250$. The comparator penalized optimization methods are examined at the same dimension and a variety of penalty values. We note the challenge in selection of the penalty $\lambda$ and emphasize that this value must be chosen prior to analysis and not fine tuned after the fact to produce a desired (already known) result. Results for the fabricated, naive scenario are compiled in Figure 4, and we demonstrate that the permutation based methodology extracts the correct relationship between the jaguar mating call sound fragment and the whole-cell recording response where the penalized optimization approaches struggle.
Figure 4: Estimated Granger causal relationship between each acoustic stimulus and primary auditory cortex response to the jaguar mating call sound fragment (NPGC quantiles $\hat{Q}_{M}=0.0025,1.000,0.4425,$ from top to bottom). The NPGC method recovers the correct causal structure, while the lasso-based cMLP methods fail to isolate the jaguar stimulus.
6 Conclusion
The NPGC methodology illustrates the use of the permutation-based shift from in-sample to out-of-sample testing. Permutation tests are used widely in literature, and Nauta et al. (2019) implements a decision rule similar to NPGC. We explicitly define Granger causality in this framework, and provide a theoretical analysis of the corresponding variance estimates and decision rule. This shift allows for control in identifying useful functional relationships when overfitting from an artificial network becomes a concern, and removes the burden of specifying regularization parameters that have a major impact on the observed outcome. Once a Granger causal outcome has been determined, penalized optimization techniques provide a practical screening method for identifying the prominent relationships, but they do not effectively estimate the global presence or absence of functional connectivity.
We show the ability of the permutation framework to detect the presence of functional connectivity, and to provide a safeguard against misidentification of data-specific noise as functional dependence when the artificial network is overfit. Alternative formulations of the featurizing function $\Psi$ in Equation 4 can utilize the same methodological structure and retain the theoretical guarantees provided the transformation meets the required conditions listed at the beginning of Section 3 and the construction allows for row-wise permutation without breaking the dependence structure across variables in $\mathbf{X}$. Misspecification of the dimension of the feature space should only produce additional false negative results as long as $N$ is less than the number of data points used in model estimation. We suggest using a large enough dimension to capture any potential nonlinear dependence structure that may exist in the dataset. We also note that the NPGC method exhibits slight undercoverage for a chosen level of test, but argue that minor deviations in this realm are not of great concern if the result is correctly interpreted as a potential functional connection and not an outright causal effect. The permutation method circumvents the need for a penalty parameter selection, and provides ease of extension to multiple testing problems and control of family-wise error rate.
We caution against potential misuses of Granger causality including, but not limited to, repeated application to subsets of a selected covariate group without adequate Type 1 error correction, disregard for the conditional nature of the inferential conclusion, attempts at individual rather than collective covariate inference (unless specification of the model is complete and exact), inclusion of a covariate without careful scientific or logical reasoning, and its use as an outright mechanism for identifying causal relationships rather than predictive links for future study. We also recommend prudent selection of the length of lag response included in the covariate matrix and recognize that two values may produce different results. In nonlinear processes, selection of the relevant history of the response for inclusion in the model remains an area of future study.
SUPPLEMENTARY MATERIAL
Nonlinear Permuted Granger Causality
Additional Tables & Figures
: Tables and figures of simulation and application results not presented in the main paper. (PDF)
Algorithms & Pseudocode:
Presentation of algorithms mentioned in Section 2 in the form of pseudocode. See main text for explanation. (PDF)
Theory Supplement:
Conditions for theoretical results and proofs for theorems presented in the paper. (PDF)
Simulation data:
Simulated data used to assess performance of methods available at
https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.RData, .npy, and .csv files)
Simulation code:
Code used to generate results and figures discussed in paper available at https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.R, .cpp, and.py files)
Application data:
Data used for applications Section 5 available at
https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.mat file)
Application code:
Code used to assess the methods in the application study available at https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.R, .py, and .m files)
7 Additional Tables & Figures
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 0.895 0.795 0.110 0.435 0.865 0.985
$T=500$ 0.975 0.805 0.130 0.505 0.910 1.000
$T=1000$ 0.990 0.805 0.090 0.420 0.915 1.000
(a) TAR(2) response with zero Granger causal variables included in additional set $\mathbf{Z}$.
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 0.935 0.945 0.200 0.675 0.640 0.990
$T=500$ 0.955 0.900 0.135 0.610 0.775 0.995
$T=1000$ 0.980 0.895 0.155 0.580 0.785 1.000
(b) TAR(2) response with two Granger causal variables included in additional set $\mathbf{Z}$.
Table 5: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for TAR(2) simulations.
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 0.955 0.705 0.220 0.435 0.050 1.000
$T=500$ 0.980 0.650 0.115 0.380 0.015 1.000
$T=1000$ 0.985 0.670 0.090 0.360 0.040 1.000
(a) Lorenz-96 response with zero Granger causal variables included in additional set $\mathbf{Z}$.
NPGC cMLP-GL cMLP-GSGL cMLP-H R/U GNS
$T=250$ 1.000 0.825 0.245 0.530 0.000 1.000
$T=500$ 0.995 0.805 0.205 0.505 0.000 1.000
$T=1000$ 1.000 0.760 0.165 0.465 0.000 1.000
(b) Lorenz-96 response with two Granger causal variables included in additional set $\mathbf{Z}$.
Table 6: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for Lorenz-96 simulations.
Figure 5: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations with zero (top) and two (bottom) Granger causal variables included in additional set $\mathbf{Z}$.
Figure 6: Type 1 Error control by reference quantile $\alpha$ for Lorenz-96 simulations with zero (top) and two (bottom) Granger causal variables included in additional set $\mathbf{Z}$.
Figure 7: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations without Lorenz-96 group included to demonstrate adherence to theoretical results. Zero Granger causal variables (left) and two Granger causal variables (right) included in the additional set $\mathbf{Z}$. All data sets are $T=1000$.
8 Algorithms for Methodology
Algorithm 1 Nonlinear Permuted Granger Causality
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # permutations $M$; # random initializations $\mathscr{R}$; feature space dimension $N$; # cross validation folds $K$
Outputs: $\hat{Q}_{M}$; $\hat{\vartheta}_{m}$ for each permutation $m=1,\ldots,M$
1:Generate permutations $\tilde{\mathbf{X}}_{m}=\bm{\Pi}_{m}\mathbf{X}$ for $m=1,\ldots,M$ with $\bm{\Pi}_{1}=\mathbf{I}$
2:Initialize $\mathbf{W}_{r}\in\mathbb{R}^{(1+\gamma d+q+p)\times N}$ where each element $w_{r,ij}\sim\mathcal{N}\left(0,1\right)$ for all $\mathscr{R}$
3:for $m$ in $1:M$ do
4:  for $\omega$ in $1:\varphi$ do
5:   for $r$ in $1:\mathscr{R}$ do
6:     $\mathbf{H}_{\omega,m,r}\leftarrow\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{m,\omega}\right]\mathbf{W}_{r}\right)$
7:     for $k$ in $1:K$ do
8:      $\mathbf{R}_{m,\omega,r,k}\leftarrow\mathbf{H}_{m,\omega,r,k}\left(\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right)^{-1}\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{Y}_{\omega,-k}-\mathbf{Y}_{\omega,k}$
9:     end for
10:   end for
11:  end for
12:  $\hat{\vartheta}_{m}\leftarrow\left(\varphi\mathscr{R}K\right)^{-1}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}T_{k}^{-1}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)$
13:end for
[…]
Automated selection of the feature space dimension in Algorithm 2 uses cross validation sets $1,\ldots,k^{*}$ of the $K$ total to obtain $N$. Implementation of Algorithm 2 requires alteration of Algorithm 1 to only use test sets $k=k^{*}+1,\ldots,K$.
Algorithm 2 Automated Feature Space Dimension Selection
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # random initializations $\mathscr{R}$; # cross validation folds $K$
Outputs: feature space dimension NN
1:Remove set $k^{*}+1,\ldots,K$ from $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ leaving only folds $1,\ldots,k^{*}$
[…]
3:for $n$ in $10:10:N_{\max}$ do
4:  Initialize $\mathbf{W}_{r}\in\mathbb{R}^{\left(1+\gamma d+q+p\right)\times n}$ where each element $w_{r,ij}\sim\mathcal{N}\left(0,1\right)$ for all $\mathscr{R}$
5:  for $\omega$ in $1:\varphi$ do
6:   for $r$ in $1:\mathscr{R}$ do
7:     $\mathbf{H}_{n,\omega,r}\leftarrow\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\mathbf{X}_{\omega}\right]\mathbf{W}_{r}\right)$
8:     for $k$ in $1:k^{*}$ do
9:      $\mathbf{R}_{n,\omega,r,k}\leftarrow\mathbf{H}_{n,\omega,r,k}\left(\mathbf{H}_{n,\omega,r,-k}^{\prime}\mathbf{H}_{n,\omega,r,-k}\right)^{-1}\mathbf{H}_{n,\omega,r,-k}^{\prime}\mathbf{Y}_{\omega,-k}-\mathbf{Y}_{\omega,k}$
10:     end for
11:   end for
12:  end for
13:end for
[…]
9 Theory Supplement
We begin with the additional conditions for theoretical results presented in Section 3.
9.1 Conditions for Theoretical Results
Define the functional transformation $\Psi$ in Equation 4 with element-wise nonlinear activation function $g=\tanh$. We derive the results for a generic $g$ under the constraints of Condition 4.
Condition 4.
The nonlinear function activation function $g$ in Equation 4 is bounded, $g:\mathbb{R}\rightarrow\left[a,b\right]$ for some $a<b$, such that for any $x\in\mathbb{R}$, $|g(x)|\leq G=\max\left\{|a|,|b|\right\}$ and $G<\infty$.
Condition 5.
Let $\mathcal{W}$ be the space of all element-wise random initializations, $w_{ij}\sim\mathcal{N}(0,1)$, such that for all $\mathbf{W}_{r}\in\mathcal{W}$, the matrix $\mathbf{W}_{r}$ is full rank and $\Psi$ generates a feature matrix that is full rank with a finite condition number. For permutation $m$, realization $\omega$, and random initialization $r$,
$\text{rank}\left(\mathbf{H}_{m,\omega,r}\right)$ $=N$ (25)
$\text{and}\qquad\kappa(\mathbf{H}_{m,\omega,r})$ $=\sigma_{1}(\mathbf{H}_{m,\omega,r})/\sigma_{N}(\mathbf{H}_{m,\omega,r})\leq\kappa_{\max}<\infty.$ (26)
All initialized model matrices $\mathbf{W}_{r}\in\mathcal{W}$.
For the matrix $\mathbf{H}_{m,\omega,r}=(h_{m,\omega,r,ij})$, define $G$ as the maximal element from Condition 4 and the average squared entry as $\bar{h}_{m,\omega,r}^{2}$.
Condition 6.
The following expected values exist and are finite:
$\mathbb{E}\left[G^{-2}\bar{h}_{m,\omega,r}^{2}|\mathbf{S}_{m,\omega,r}\right]$ $=\nu^{2}<\infty,$ (27)
$\mathbb{E}\left[G^{2}\left(\bar{h}_{m,\omega,r}^{2}\right)^{-1}|\mathbf{S}_{m,\omega,r}\right]$ $=\xi^{2}<\infty,$ (28)
$\text{and}\quad\mathbb{E}\left[G^{4}\left(\bar{h}_{m,\omega,r}^{2}\right)^{-2}|\mathbf{S}_{m,\omega,r}\right]$ $=\varrho^{4}<\infty.$ (29)
Combining Conditions 4, 5, and 6, $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$. We define $f$ as the true functional relationship between the response and the unpermuted predictor matrix for all $\omega\in\Omega$, $\mathbf{Y}_{\omega}=f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)+\bm{\varepsilon}_{\omega}$, and the approximating model form for permutations $m=1,\ldots,M$ and initializations $r=1,\ldots,\mathscr{R}$.
$\mathbf{Y}_{\omega}=\mathbf{H}_{m,\omega,r}\mathbf{B}_{m,\omega,r}+\mathbf{U}_{m,\omega,r}$ (30)
Condition 7.
For every $\eta>0$, there exists a fixed $N$, where $1+\gamma d+q+p\leq N<\infty$, such that as the number of random initializations $\mathscr{R}\rightarrow\infty$,
$\sup_{\omega\in\Omega}\left\|\frac{1}{\mathscr{R}}\sum_{r=1}^{\mathscr{R}}\mathbf{H}_{1,\omega,r}\mathbf{B}_{1,\omega,r}-f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)\right\|<\eta,$ (31)
where $\mathbf{B}_{1,\omega,r}$ is the true coefficient matrix of feature space $r$ for the unpermuted covariate set, $\mathbf{H}_{1,\omega,r}=\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\mathbf{W}_{r}\right)$.
We note that Condition 7 specifically pertains to the $f$ piece of the true functional relationship; we make no assumption of the closeness of a transformed response $\mathbf{H}_{1,\omega,r}\mathbf{B}_{1,\omega,r}$ and the true values $\mathbf{Y}_{\omega}$ if the predictors themselves are not reliable and the entries of $\bm{\varepsilon}_{\omega}$ are large.
9.2 Proofs
Lemma 1.
Under the conditions listed in Section 3 and the supplement, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$.
The proof of Lemma 1 begins with the estimate for one test set with a single individual random initialization given the variation measure $\mathbf{S}_{m,\omega,r}$. We aggregate these expectations over the sets $k=1,\ldots,K$, realizations $r=1,\ldots,\mathscr{R}$ and the $\varphi\geq 1$ observed datasets in $\Omega_{\text{obs}}$.
Proof of Lemma 1.
We seek the conditional expectation $\mathbb{E}\left[\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$, and can write the unconditional expectation of the underlying parameter based on the definition in Equation 6.
$\mathbb{E}\left[\hat{\vartheta}_{m}\right]$ $=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\mathbb{E}\left[\mathbb{E}\left(\mathbb{E}\left[\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (32)
From Condition 3, we can write the distribution of the residual term $\mathbf{R}_{m,\omega,r,k}$ from that of $\mathbf{U}_{m,\omega,r}$.
$\mathbf{U}_{m,\omega,r}$ $\sim\mathcal{MN}_{T\times d}\left(\mathbf{0},\mathbf{I}_{T},\mathbf{S}_{m,\omega,r}\right)$ (33)
$\mathbf{R}_{m,\omega,r,k}$ $=\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{Y}_{\omega,-k}-\mathbf{Y}_{\omega,k}$ (34)
$\mathbf{R}_{m,\omega,r,k}$ $\sim\mathcal{MN}_{T_{k}\times d}\left(\mathbf{0},\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}+\mathbf{I}_{T_{k}},\mathbf{S}_{m,\omega,r}\right)$ (35)
Define $\bm{\Phi}_{m,\omega,r,k}=\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}$, and further define $\mathbf{V}_{m,\omega,r,k}\sim\mathcal{MN}_{T_{k}\times d}(\mathbf{0},\mathbf{I}_{T_{k}},\mathbf{I}_{d})$ such that we can write $\mathbf{R}_{m,\omega,r,k}=\left(\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right)^{1/2}\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}^{1/2}$.
$\mathbb{E}\left[\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$
$\hskip 72.26999pt=\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}^{1/2}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}^{1/2}\right)|\mathbf{S}_{m,\omega,r}\right]$ (36)
$\hskip 72.26999pt=\mathbb{E}\left[\text{tr}\left(\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ (37)
$\hskip 72.26999pt=\text{tr}\left(\mathbb{E}\left[\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]\right)$ (38)
$\hskip 72.26999pt=\text{tr}\left(\mathbf{S}_{m,\omega,r}\mathbb{E}\left[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]\right)$ (39)
We can write
$\mathbb{E}\left[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]$
$\hskip 72.26999pt=\mathbb{E}\left[\mathbb{E}\left(\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\bm{\Phi}_{m,\omega,r,k},\mathbf{S}_{m,\omega,r}\right)\big|\mathbf{S}_{m,\omega,r}\right],$ (40)
and from Gupta and Nagar (2018),
$\mathbb{E}\left(\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\bm{\Phi}_{m,\omega,r,k},\mathbf{S}_{m,\omega,r}\right)]$
$\hskip 72.26999pt=\text{tr}\left(\mathbf{I}_{T_{k}}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\right)\mathbf{I}_{d}$ (41)
$\hskip 72.26999pt=\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right)\mathbf{I}_{d}.$ (42)
Thus,
$\mathbb{E}\left[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]$ $=\mathbb{E}\left[\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right)\mathbf{I}_{d}|\mathbf{S}_{m,\omega,r}\right]$ (43)
$=\mathbb{E}\left[\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]\mathbf{I}_{d}+T_{k}\mathbf{I}_{d}.$ (44)
We modify Equation 44 by exchanging $\bm{\Phi}_{m,\omega,r,k}$ for its definition as written above.
$\mathbb{E}\left[\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ $=\mathbb{E}\left[\text{tr}\left(\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}\right)|\mathbf{S}_{m,\omega,r}\right]$ (45)
$=\mathbb{E}\left[\text{tr}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ (46)
We examine the matrices $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ under Conditions 4, 5, 6, and 7. Recall $T_{k}$ is the number of rows in the test set, and $T_{-k}$ the number in the training set. We obtain a fixed and finite $N$ from Condition 7, and $\text{rank}(\mathbf{H}_{m,\omega,r,-k})=N$ under Equation 25 from Condition 5. We label the singular values of the matrices $\sigma_{i}(\mathbf{H}_{m,\omega,r,-k})>0$ and $\sigma_{i}(\mathbf{H}_{m,\omega,r,k})\geq 0$ for $i=1,\ldots,N$. For any matrix $\mathbf{P}$, $\left\|\mathbf{P}\right\|_{F}=\sqrt{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|p_{ij}|^{2}}=\sqrt{\sum_{i=1}^{\text{rank}(\mathbf{P})}\sigma_{i}(\mathbf{P})^{2}}$. From Condition 4, each entry in $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ lies in a bounded interval, $h_{m,\omega,r,-k,ij},h_{m,\omega,r,k,ij}\in[a,b]$ such that $|h_{m,\omega,r,-k,ij}|,|h_{m,\omega,r,k,ij}|\leq G$, and we use this to establish the upper bounds for the maximal singular values shown in Equations 47 and 48. To write the minimum values, we define the average squared matrix entry as in Condition 6, and note the combination of Conditions 4, 5, and 6 yields $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$.
$\sqrt{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ $\leq\sigma_{1}\left(\mathbf{H}_{m,\omega,r,-k}\right)\leq\left\|\mathbf{H}_{m,\omega,r,-k}\right\|_{F}\leq\sqrt{T_{-k}NG^{2}}$ (47)
$\sqrt{T_{k}\bar{h}^{2}_{m,\omega,r,k}}$ $\leq\sigma_{1}\left(\mathbf{H}_{m,\omega,r,k}\right)\leq\left\|\mathbf{H}_{m,\omega,r,k}\right\|_{F}\leq\sqrt{T_{k}NG^{2}}$ (48)
Under Condition 5, $\mathbf{H}_{m,\omega,r,-k}$ has a finite condition number $\kappa\left(\mathbf{H}_{m,\omega,r,-k}\right)\leq\kappa_{\max}$, and we can use the bounds in Equations 47 and 48 to obtain bounds for the minimum singular values. The rank of $\mathbf{H}_{m,\omega,r,k}$ is not necessarily $N$, as $T_{k}$ can be less than $N$, but it is at least $\min\left\{T_{k},N\right\}$.
$\frac{\sqrt{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}}{\kappa_{\max}}$ $\leq\sigma_{N}\left(\mathbf{H}_{m,\omega,r,-k}\right)\leq\sqrt{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ (49)
$0$ $\leq\sigma_{N}\left(\mathbf{H}_{m,\omega,r,k}\right)\leq\sqrt{T_{k}\bar{h}^{2}_{m,\omega,r,k}}$ (50)
We return to the $\text{tr}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$ piece of Equation 46. Applying the above results, we can write $\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\succ 0$, $\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\succeq 0$, and the following bounds on the extreme eigenvalues.
$\frac{1}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ $\leq\lambda_{1}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\leq\frac{\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ (51)
$\frac{1}{T_{-k}NG^{2}}$ $\leq\lambda_{N}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\leq\frac{1}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ (52)
$T_{k}\bar{h}^{2}_{m,\omega,r,k}$ $\leq\lambda_{1}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)\leq T_{k}NG^{2}$ (53)
$0$ $\leq\lambda_{N}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)\leq T_{k}\bar{h}^{2}_{m,\omega,r,k}$ (54)
We can bound the trace of the product with the eigenvalues using Von Neumann’s trace inequality.
$\text{tr}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$
$\hskip 72.26999pt\leq\sum_{i=1}^{N}\lambda_{i}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\lambda_{i}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$ (55)
$\hskip 72.26999pt\leq N\lambda_{1}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\lambda_{1}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$ (56)
$\hskip 72.26999pt\leq\frac{T_{k}N^{2}G^{2}\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}$ (57)
The upper bound in Equation 57 is a positive, finite constant. Returning to Equation 46,
$\mathbb{E}\left[\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ $\leq\mathbb{E}\left[\frac{T_{k}N^{2}G^{2}\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}\big|\mathbf{S}_{m,\omega,r}\right]$ (58)
$\leq\frac{T_{k}}{T_{-k}}\left(NG\kappa_{\max}\right)^{2}\mathbb{E}\left[\left(\bar{h}_{m,\omega,r,-k}^{2}\right)^{-1}|\mathbf{S}_{m,\omega,r}\right]$ (59)
$\leq\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}.$ (60)
Using this result in Equation 44,
$\mathbb{E}\left[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]$ $\leq\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}\mathbf{I}_{d}+T_{k}\mathbf{I}_{d},$ (61)
and in Equation 39,
$\mathbb{E}\left[\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ $\leq\text{tr}\left(\mathbf{S}_{m,\omega,r}\left[\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}\mathbf{I}_{d}+T_{k}\mathbf{I}_{d}\right]\right)$ (62)
$\leq\left[\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+T_{k}\right]\text{tr}\left(\mathbf{S}_{m,\omega,r}\right).$ (63)
We now require a lower bound for the expectation. We examine the trace inequality
$\text{tr}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$
$\hskip 72.26999pt\geq\sum_{i=1}^{N}\lambda_{i}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\lambda_{N-i+1}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$ (64)
$\hskip 72.26999pt\geq\lambda_{N}\left(\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\right)\lambda_{1}\left(\mathbf{H}_{m,\omega,r,k}^{\prime}\mathbf{H}_{m,\omega,r,k}\right)$ (65)
$\hskip 72.26999pt\geq\frac{T_{k}\bar{h}^{2}_{m,\omega,r,k}}{T_{-k}NG^{2}}.$ (66)
Again returning to Equation 46,
$\mathbb{E}\left[\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ $\geq\mathbb{E}\left[\frac{T_{k}\bar{h}^{2}_{m,\omega,r,k}}{T_{-k}NG^{2}}\big|\mathbf{S}_{m,\omega,r}\right]$ (67)
$\geq\frac{T_{k}}{T_{-k}NG^{2}}\mathbb{E}\left[\bar{h}_{m,\omega,r,k}^{2}|\mathbf{S}_{m,\omega,r}\right]$ (68)
$\geq\frac{T_{k}\nu^{2}}{T_{-k}N},$ (69)
and using the result in Equation 44,
$\mathbb{E}\left[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\right]$ $\geq\frac{T_{k}\nu^{2}}{T_{-k}N}\mathbf{I}_{d}+T_{k}\mathbf{I}_{d}.$ (70)
This yields the lower bound,
$\mathbb{E}\left[\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$ $\geq\text{tr}\left(\mathbf{S}_{m,\omega,r}\left[\frac{T_{k}\nu^{2}}{T_{-k}N}\mathbf{I}_{d}+T_{k}\mathbf{I}_{d}\right]\right)$ (71)
$\geq\left[\frac{T_{k}\nu^{2}}{T_{-k}N}+T_{k}\right]\text{tr}\left(\mathbf{S}_{m,\omega,r}\right).$ (72)
We return to the expectation equation given in Equation 32 with the two bounds from Equations 63 and 72.
$\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\mathbb{E}\left[\mathbb{E}\left(\left[\frac{T_{k}\nu^{2}}{T_{-k}N}+T_{k}\right]\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right)\right]\leq\mathbb{E}\left[\hat{\vartheta}_{m}\right]$
$\hskip 36.135pt\leq\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\mathbb{E}\left[\mathbb{E}\left(\left[\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+T_{k}\right]\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right)\right]$ (73)
$\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\mathbb{E}\left[\left(\frac{T_{k}\nu^{2}}{T_{-k}N}+T_{k}\right)\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]\leq\mathbb{E}\left[\hat{\vartheta}_{m}\right]$
$\hskip 36.135pt\leq\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\mathbb{E}\left[\left(\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+T_{k}\right)\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right]$ (74)
$\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\left[\frac{\nu^{2}}{T_{-k}N}+1\right]\vartheta_{m}\leq\mathbb{E}\left[\hat{\vartheta}_{m}\right]$
$\hskip 36.135pt\leq\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\left[\frac{1}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+1\right]\vartheta_{m}$ (75)
In the limit as $T\rightarrow\infty$, with a fixed test set size $T_{k}<\infty$, $T_{-k}=T-T_{k}\rightarrow\infty$.
$\lim_{T\rightarrow\infty}\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\left[\frac{\nu^{2}}{T_{-k}N}+1\right]\vartheta_{m}$
$\hskip 72.26999pt=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\lim_{T_{-k}\rightarrow\infty}\left[\frac{\nu^{2}}{T_{-k}N}+1\right]\vartheta_{m}=\vartheta_{m}$ (76)
$\lim_{T\rightarrow\infty}\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\left[\frac{1}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+1\right]\vartheta_{m}$
$\hskip 72.26999pt=\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\lim_{T_{-k}\rightarrow\infty}\left[\frac{1}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+1\right]\vartheta_{m}=\vartheta_{m}$ (77)
Thus, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$. ∎
Lemma 2.
Under the conditions listed in Section 3 and the supplement, $\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)=\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$.
For use in the proof of Lemma 2, we first state and prove Remark 1, and then proceed like in the proof of Lemma 1.
Remark 1.
For any square, positive semidefinite matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$, $\text{tr}^{2}\left(\mathbf{P}\right)\leq(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$.
Proof of Remark 1.
Let $\lambda_{1}\left(\mathbf{P}\right),\lambda_{2}\left(\mathbf{P}\right),\ldots,\lambda_{n}\left(\mathbf{P}\right)\geq 0$ be the eigenvalues of $\mathbf{P}\in\mathbb{R}^{n\times n}$. For any $n\geq 1$ we can write,
$\text{tr}^{2}\left(\mathbf{P}\right)$ $=\left(\sum_{i=1}^{n}\lambda_{i}(\mathbf{P})\right)^{2}$ (78)
$=\sum_{i=1}^{n}\lambda_{i}\left(\mathbf{P}\right)^{2}+2\sum_{i=1}^{n}\sum_{j=1}^{i-1}\lambda_{i}\left(\mathbf{P}\right)\lambda_{j}\left(\mathbf{P}\right)$ (79)
$\leq\sum_{i=1}^{n}\lambda_{i}\left(\mathbf{P}\right)^{2}+n(n-1)\sum_{i=1}^{n}\lambda_{i}\left(\mathbf{P}\right)^{2}$ (80)
$\leq\left(n^{2}-n+1\right)\sum_{i=1}^{n}\lambda_{i}\left(\mathbf{P}\right)^{2}$ (81)
with the last line equal to $(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$. Thus, $\text{tr}^{2}\left(\mathbf{P}\right)\leq(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$. ∎
Proof of Lemma 2.
We isolate the quantity $\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)$ like in the proof of Lemma 1 and examine the conditional variance $\text{Var}\left(\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$. The unconditional variance follows from the law of total variance and from the definition in Equation 6.
$\text{Var}\left(\hat{\vartheta}_{m}\right)$ $=\mathbb{E}\left[\mathbb{E}\left(\text{Var}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$\qquad+\mathbb{E}\left[\text{Var}\left(\mathbb{E}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$\qquad+\text{Var}\left(\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left[\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$ (82)
We first examine $\text{Var}\left(\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$.
$\text{Var}\left(\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)=\mathbb{E}\left[\text{tr}^{2}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$
$\hskip 216.81pt-\mathbb{E}\left[\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]^{2}$ (83)
To establish an bounds on the variance, we seek both lower and upper bounds for the terms in Equation 83. The bounds for the second term can be directly obtained from the derivation in the Proof of Lemma 1.
$\left[\frac{T_{k}\nu^{2}}{T_{-k}N}+T_{k}\right]^{2}\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)\leq\mathbb{E}\left[\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]^{2}$
$\hskip 108.405pt\leq\left[\frac{T_{k}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+T_{k}\right]^{2}\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)$ (84)
$\left[\frac{T_{k}^{2}\nu^{4}}{T_{-k}^{2}N^{2}}+2\frac{T_{k}^{2}\nu^{2}}{T_{-k}N}+T_{k}^{2}\right]\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)\leq\mathbb{E}\left[\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]^{2}$
$\hskip 108.405pt\leq\left[\frac{T_{k}^{2}}{T_{-k}^{2}}\left(N\xi\kappa_{\max}\right)^{4}+2\frac{T_{k}^{2}}{T_{-k}}\left(N\xi\kappa_{\max}\right)^{2}+T_{k}^{2}\right]\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)$ (85)
For the first term in Equation 83, we follow the strategy of the proof of Lemma 1. The matrix in the trace of Equation 86 is square and positive semidefinite, and we apply the result of Remark 1.
$\mathbb{E}\left[\text{tr}^{2}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$
$\hskip 36.135pt=\mathbb{E}\left[\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}^{1/2}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}^{1/2}\right)|\mathbf{S}_{m,\omega,r}\right]$ (86)
$\hskip 36.135pt\leq(d^{2}-d+1)\;\mathbb{E}\bigg[\text{tr}\bigg(\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}$
$\hskip 180.67499pt\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\bigg)|\mathbf{S}_{m,\omega,r}\bigg]$ (87)
$\hskip 36.135pt\leq(d^{2}-d+1)\;\text{tr}\bigg(\mathbf{S}_{m,\omega,r}\mathbb{E}\bigg[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}$
$\hskip 180.67499pt\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\bigg]\bigg)$ (88)
We can write the expectation as
$\mathbb{E}\bigg[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\bigg]$
$\hskip 36.135pt=\mathbb{E}\bigg[\mathbb{E}\bigg(\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}$
$\hskip 108.405pt\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\bm{\Phi}_{m,\omega,r,k},\mathbf{S}_{m,\omega,r}\bigg)|\mathbf{S}_{m,\omega,r}\bigg]$ (89)
and obtain the inner piece from Theorem 2.3.8 (v) of Gupta and Nagar (2018).
$\mathbb{E}\bigg(\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\bm{\Phi}_{m,\omega,r,k},\mathbf{S}_{m,\omega,r}\bigg)$
$\hskip 36.135pt=\text{tr}\left(\mathbf{I}_{T_{k}}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{I}_{T_{k}}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\mathbf{I}_{d}\right)\mathbf{I}_{d}$
$\hskip 72.26999pt+\text{tr}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{I}_{T_{k}}\right)\text{tr}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{I}_{T_{k}}\right)\mathbf{I}_{d}\mathbf{S}_{m,\omega,r}\mathbf{I}_{d}$
$\hskip 72.26999pt+\text{tr}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{I}_{T_{k}}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{I}_{T_{k}}\right)\mathbf{I}_{d}\mathbf{S}_{m,\omega,r}\mathbf{I}_{d}$ (90)
$\hskip 36.135pt=\text{tr}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]^{2}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}$
$\hskip 72.26999pt+\text{tr}^{2}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+\text{tr}\left(\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]^{2}\right)\mathbf{S}_{m,\omega,r}$ (91)
$\hskip 36.135pt=\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}^{2}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}^{2}\right)\mathbf{S}_{m,\omega,r}+\text{tr}^{2}\left(\bm{\Phi}_{m,\omega,r,k}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+2\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+2\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)\mathbf{S}_{m,\omega,r}+2T_{k}\text{tr}\left(\bm{\Phi}_{m,\omega,r,k}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+T_{k}\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+T_{k}\mathbf{S}_{m,\omega,r}+T_{k}^{2}\mathbf{S}_{m,\omega,r}$ (92)
Using Von Neumann’s trace inequality and the upper bound on the largest eigenvalue from Equation 57,
$\text{tr}\left(\bm{\Phi}^{2}_{m,\omega,r,k}\right)$ $\leq\sum_{i=1}^{N}\lambda_{i}\left(\bm{\Phi}_{m,\omega,r,k}\right)^{2}$ (93)
$\leq N\lambda_{1}\left(\bm{\Phi}_{m,\omega,r,k}\right)^{2}$ (94)
$\leq\frac{T_{k}^{2}N^{3}G^{4}\kappa_{\max}^{4}}{T_{-k}^{2}\left(\bar{h}^{2}_{m,\omega,r,-k}\right)^{2}},$ (95)
and we input the result to Equation 92.
$\mathbb{E}\bigg(\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\bm{\Phi}_{m,\omega,r,k},\mathbf{S}_{m,\omega,r}\bigg)$
$\hskip 36.135pt\leq\left(\frac{T_{k}^{2}N^{3}G^{4}\kappa_{\max}^{4}}{T_{-k}^{2}\left(\bar{h}^{2}_{m,\omega,r,-k}\right)^{2}}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+\left(\frac{T_{k}^{2}N^{3}G^{4}\kappa_{\max}^{4}}{T_{-k}^{2}\left(\bar{h}^{2}_{m,\omega,r,-k}\right)^{2}}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+\left(\frac{T_{k}^{2}N^{4}G^{4}\kappa_{\max}^{4}}{T_{-k}^{2}\left(\bar{h}_{m,\omega,r,-k}^{2}\right)^{2}}\right)\mathbf{S}_{m,\omega,r}+2\left(\frac{T_{k}N^{2}G^{2}\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}$
$\hskip 72.26999pt+2\left(\frac{T_{k}N^{2}G^{2}\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}\right)\mathbf{S}_{m,\omega,r}+2\left(\frac{T_{k}^{2}N^{2}G^{2}\kappa_{\max}^{2}}{T_{-k}\bar{h}_{m,\omega,r,-k}^{2}}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+T_{k}\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+T_{k}\mathbf{S}_{m,\omega,r}+T_{k}^{2}\mathbf{S}_{m,\omega,r}$ (96)
With the expected values from Condition 6, we can simplify the form to that shown in Equation 97.
$\mathbb{E}\bigg[\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}\mathbf{V}^{\prime}_{m,\omega,r,k}\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\mathbf{V}_{m,\omega,r,k}|\mathbf{S}_{m,\omega,r}\bigg]$
$\hskip 36.135pt\leq\left(\frac{T_{k}^{2}N^{3}\varrho^{4}\kappa_{\max}^{4}}{T_{-k}^{2}}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+\left(\frac{T_{k}^{2}N^{3}\varrho^{4}\kappa_{\max}^{4}}{T_{-k}^{2}}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+\left(\frac{T_{k}^{2}N^{4}\varrho^{4}\kappa_{\max}^{4}}{T_{-k}^{2}}\right)\mathbf{S}_{m,\omega,r}+2\left(\frac{T_{k}N^{2}\xi^{2}\kappa_{\max}^{2}}{T_{-k}}\right)\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}$
$\hskip 72.26999pt+2\left(\frac{T_{k}N^{2}\xi^{2}\kappa_{\max}^{2}}{T_{-k}}\right)\mathbf{S}_{m,\omega,r}+2\left(\frac{T_{k}^{2}N^{2}\xi^{2}\kappa_{\max}^{2}}{T_{-k}}\right)\mathbf{S}_{m,\omega,r}$
$\hskip 72.26999pt+T_{k}\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)\mathbf{I}_{d}+T_{k}\mathbf{S}_{m,\omega,r}+T_{k}^{2}\mathbf{S}_{m,\omega,r}$ (97)
We return to Equation 88, and note that for any square, positive semidefinite matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$, $\text{tr}\left(\mathbf{P}^{2}\right)\leq\text{tr}^{2}\left(\mathbf{P}\right)$.
$\mathbb{E}\left[\text{tr}^{2}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]$
$\hskip 36.135pt\leq(d^{2}-d+1)\;\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)\bigg[3\left(\frac{T_{k}^{2}N^{4}\varrho^{4}\kappa_{\max}^{4}}{T_{-k}^{2}}\right)+4\left(\frac{T_{k}N^{2}\xi^{2}\kappa_{\max}^{2}}{T_{-k}}\right)$
$\hskip 108.405pt+2\left(\frac{T_{k}^{2}N^{2}\xi^{2}\kappa_{\max}^{2}}{T_{-k}}\right)+2T_{k}+T_{k}^{2}\bigg]$ (98)
We can write the following inequality for $\text{Var}\left(T_{k}^{-1}\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$.
$\text{Var}\left(T_{k}^{-1}\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$
$\hskip 36.135pt\leq\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)\bigg(\frac{1}{T_{-k}^{2}}\left[\frac{3(d^{2}-d+1)N^{6}\varrho^{4}\kappa_{\max}^{4}-\nu^{4}}{N^{2}}\right]$
$\hskip 72.26999pt+\frac{1}{T_{k}T_{-k}}\left[4(d^{2}-d+1)N^{2}\xi^{2}\kappa_{\max}^{2}\right]$
$\hskip 72.26999pt+\frac{1}{T_{-k}}\left[\frac{2(d^{2}-d+1)N^{3}\xi^{2}\kappa_{\max}^{2}-2\nu^{2}}{N}\right]$
$\hskip 72.26999pt+(d^{2}-d)+\frac{2}{T_{K}}(d^{2}-d+1)\bigg)$ (99)
Taking the limit as $T\rightarrow\infty$, we obtain an upper bound for the variance of the estimate given all individual covariance matrices for one test set with each initialization and potential realization of the data. As the number of time points gets large with a fixed $T_{k}<\infty$, the size of each training set $T_{-k}$ tends to infinity.
$\lim_{T\rightarrow\infty}\text{Var}\left(T_{k}^{-1}\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$ $\leq\left[(d^{2}-d)+\frac{2}{T_{k}}(d^{2}-d+1)\right]\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)$ (100)
Returning to the total variance in Equation 82, with the limit in Equation 100, we can apply the Dominated Convergence Theorem to establish the limiting value for $\text{Var}\left(\hat{\vartheta}_{m}\right)$. We also use the results in Equations 63 and 72 to establish the limiting expectation $\lim_{T\rightarrow\infty}\mathbb{E}\left[T_{k}^{-1}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]=\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)$.
The first term in Equation 82 has an inner term of the variance of the variation parameter given the initialization specific matrix $\mathbf{S}_{m,\omega,r}$ and the realization specific matrix $\bm{\Sigma}_{m,\omega}$. The only remaining source of variation arises from labelling the training and test sets $k=1,\ldots,K$. With random assignment, and fully explained temporal dependence as in Condition 3, we treat these splits as uncorrelated draws, and the variance can move inside the summation.
$\mathbb{E}\left[\mathbb{E}\left(\text{Var}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$\hskip 36.135pt=\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\text{Var}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (101)
The inner variance converges pointwise and is dominated by the integrable quantity
$\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)\bigg[3(d^{2}-d+1)N^{6}\varrho^{4}\kappa_{\max}^{4}+6(d^{2}-d+1)N^{3}\xi^{2}\kappa_{\max}^{2}+3d^{2}+1\bigg],$ (102)
as a simplification of the form shown in Equation 99. The expectation of the inner term converges pointwise, where
$\lim_{T\rightarrow\infty}\mathbb{E}\left[\text{Var}\left(T_{k}^{-1}\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]$
$\hskip 72.26999pt\leq\left[(d^{2}-d)+\frac{2}{T_{k}}(d^{2}-d+1)\right]\text{tr}^{2}\left(\bm{\Sigma}_{m,\omega}\right),$ (103)
and is dominated by the similar integrable quantity
$\text{tr}^{2}\left(\bm{\Sigma}_{m,\omega}\right)\bigg[3(d^{2}-d+1)N^{6}\varrho^{4}\kappa_{\max}^{4}+6(d^{2}-d+1)N^{3}\xi^{2}\kappa_{\max}^{2}+3d^{2}+1\bigg].$ (104)
We again assume a fixed $T_{k}<\infty$, and $T\rightarrow\infty$ implies $T_{-k}\rightarrow\infty$. The limit also implies $K\rightarrow\infty$ for a fixed $T_{k}$. When applying the Dominated Convergence Theorem twice,
$\lim_{T\rightarrow\infty}\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\text{Var}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$=\mathbb{E}\left[\lim_{T\rightarrow\infty}\mathbb{E}\left(\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\text{Var}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (105)
$=\mathbb{E}\left[\mathbb{E}\left(\lim_{T_{-k},K\rightarrow\infty}\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\text{Var}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (106)
$\leq\mathbb{E}\left[\mathbb{E}\left(\lim_{K\rightarrow\infty}\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\left[(d^{2}-d)+\frac{2}{T_{k}}(d^{2}-d+1)\right]\text{tr}^{2}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right)\right].$ (107)
The inner piece tends to zero as $K$ gets large, and
$\lim_{T\rightarrow\infty}\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi^{2}\mathscr{R}^{2}K^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\text{Var}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]=0.$ (108)
For the second term in Equation 82, the inner expectation converges pointwise to $\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)$ and is dominated by the integrable term $\left[\left(N\xi\kappa_{\max}\right)^{2}+1\right]\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)$. The middle expectation converges pointwise to $\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ and is dominated by the similar term $\left[\left(N\xi\kappa_{\max}\right)^{2}+1\right]\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$. We can take the limit inside the function $h(x)=x^{2}$ as it is continuous on the full domain $x\in\mathbb{R}$ (i.e., inside the variance term).
$\lim_{T\rightarrow\infty}\mathbb{E}\left[\text{Var}\left(\mathbb{E}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$=\mathbb{E}\left[\lim_{T\rightarrow\infty}\text{Var}\left(\mathbb{E}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (109)
$=\mathbb{E}\left[\text{Var}\left(\lim_{T\rightarrow\infty}\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\mathbb{E}\left[\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$ (110)
$=\mathbb{E}\left[\text{Var}\left(\frac{1}{\varphi\mathscr{R}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right)\right]$ (111)
Given the data realized covariance matrix $\bm{\Sigma}_{m,\omega}$, each initialization $\mathbf{S}_{m,\omega,r}$ is an uncorrelated draw, and we can distribute the variance term like above. From Condition 2, we extract the limiting value.
$\lim_{T\rightarrow\infty}\mathbb{E}\left[\text{Var}\left(\mathbb{E}\left[\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)|\mathbf{S}_{m,\omega,r}\right]|\bm{\Sigma}_{m,\omega}\right)\right]$
$=\mathbb{E}\left[\frac{1}{\varphi^{2}\mathscr{R}^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\text{Var}\left(\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right)\right]$ (112)
$=\mathbb{E}\left[\frac{1}{\varphi^{2}\mathscr{R}^{2}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\tau_{r}^{2}\right]=\frac{\tau_{r}^{2}}{\varphi\mathscr{R}}$ (113)
For the third term in Equation 82, we note the same conditions as before, and proceed like above.
$\lim_{T\rightarrow\infty}\text{Var}\left(\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left[\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$
$=\text{Var}\left(\lim_{T\rightarrow\infty}\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left[\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$ (114)
$=\text{Var}\left(\mathbb{E}\left[\lim_{T\rightarrow\infty}\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\mathbb{E}\left(\frac{1}{T_{k}}\text{tr}\left[\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$ (115)
$=\text{Var}\left(\mathbb{E}\left[\frac{1}{\varphi\mathscr{R}}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\text{tr}\left(\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$ (116)
$=\text{Var}\left(\frac{1}{\varphi}\sum_{\omega=1}^{\varphi}\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\right)$ (117)
Each realization of the data is independent, and we can take the variance inside the summation.
$\lim_{T\rightarrow\infty}\text{Var}\left(\mathbb{E}\left[\mathbb{E}\left(\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\text{tr}\left[\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)|\bm{\Sigma}_{m,\omega}\right]\right)$
$=\frac{1}{\varphi^{2}}\sum_{\omega=1}^{\varphi}\text{Var}\left(\text{tr}\left[\bm{\Sigma}_{m,\omega}\right]\right)=\frac{\tau_{\omega}^{2}}{\varphi}$ (118)
Thus, we can write the limiting variance as,
$\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)$ $=\frac{\tau_{\omega}^{2}}{\varphi}+\frac{\tau_{r}^{2}}{\varphi\mathscr{R}}.$ (119)
The proof of Theorem 1 follows directly from the results of Lemmas 1 and 2.
Proof of Theorem 1.
From Lemma 1, we establish that $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$. Turning to the result of Lemma 2,
$\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)$ $=\frac{\tau_{\omega}^{2}}{\varphi}+\frac{\tau_{r}^{2}}{\varphi\mathscr{R}}$ (120)
$\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)$ $=\lim_{\mathscr{R}\rightarrow\infty}\frac{\tau_{\omega}^{2}}{\varphi}+\frac{\tau_{r}^{2}}{\varphi\mathscr{R}}=\frac{\tau_{\omega}^{2}}{\varphi}.$ (121)
When we observe the data without error ($\tau_{\omega}^{2}=0$), the limiting variance is zero, and the estimate is consistent.
In the alternative scenario, we no longer require error free data observation.
$\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)$ $=\lim_{\varphi\rightarrow\infty}\frac{\tau_{\omega}^{2}}{\varphi}+\frac{\tau_{r}^{2}}{\varphi\mathscr{R}}=0.$ (122)
The proof of Theorem 2 is a direct result of Condition 2.
Proof of Theorem 2.
From Condition 2, we see that $\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)\overset{\textit{i.i.d.}}{\sim}\left(\vartheta_{m},\tau_{\omega}^{2}\right)$. For an individual realization, we have shown in the proofs of Lemma 1 and 2 that
$\frac{1}{\mathscr{R}K}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}T_{k}^{-1}\text{tr}\left(\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right)|\bm{\Sigma}_{m,\omega}\xrightarrow{p}\text{tr}\left(\bm{\Sigma}_{m,\omega}\right)$ (123)
as the number of observations $T\rightarrow\infty$. The variation parameter is the average of the individual realization-specific covariance matrices, and the Central Limit Theorem arises from the application of Slutsky’s Theorem. ∎
The proof of Theorem 3 begins with the empirical distribution function, and employs the result of the Glivenko-Cantelli to show convergence in distribution to a $\text{Uniform}(0,1)$ random variable.
Proof of Theorem 3.
Under the null hypothesis, $\vartheta_{1}=\cdots=\vartheta_{T!}=\vartheta$. We define the quantile estimate $\hat{Q}_{M}$, shown in Equation 9, as an evaluation of the empirical distribution function $\hat{\mathcal{H}}_{M}(s)$ at $\hat{\vartheta}_{1}$. As the number of permutations $M$ approaches the total number of possible permutations $T!$, the empirical distribution $\hat{\mathcal{H}}_{M}(s)$ trivially converges to $\hat{\mathcal{H}}(s)$ defined in Equation 7.
From the results of Lemmas 1 and 2, we know that as $T\rightarrow\infty$, $\hat{\vartheta}_{1}\xrightarrow{d}\mathcal{F}$, where the continuous distribution $\mathcal{F}$ has expectation $\vartheta$ and variance $\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$. Under the null, each realization $\hat{\vartheta}_{m}$ is independently drawn from $\mathcal{F}$. From the Glivenko-Cantelli Theorem as $T\rightarrow\infty$,
$\sup_{s}\left|\hat{\mathcal{H}}(s)-\mathcal{F}(s)\right|\xrightarrow{a.s.}0.$ (124)
We define $\mathcal{U}\sim\text{Uniform}(0,1)$ and note that if $s\sim\mathcal{G}$, we can write $s\sim\mathcal{G}^{-1}(\mathcal{U})$. We state that $\mathcal{F}\left(\lim_{T\rightarrow\infty}\hat{\vartheta}_{1}\right)$ follows the same distribution as $\mathcal{F}\left(\mathcal{F}^{-1}\left(\mathcal{U}\right)\right)\sim\mathcal{U}$.
Combining these results and Theorem 1 with the continuous mapping theorem in Equation 127 and Slutsky’s Theorem,
$\lim_{M\rightarrow T!}\hat{\mathcal{H}}_{M}\left(\hat{\vartheta}_{1}\right)$ $\rightarrow\hat{\mathcal{H}}\left(\hat{\vartheta}_{1}\right)$ (125)
$\lim_{T\rightarrow\infty}\hat{\mathcal{H}}\left(\hat{\vartheta}_{1}\right)$ $\rightarrow\mathcal{F}\left(\hat{\vartheta}_{1}\right)$ (126)
$\lim_{T\rightarrow\infty}\mathcal{F}\left(\hat{\vartheta}_{1}\right)$ $\xrightarrow{d}\mathcal{F}\left(\lim_{T\rightarrow\infty}\hat{\vartheta}_{1}\right)$ (127)
$\mathcal{F}\left(\lim_{T\rightarrow\infty}\hat{\vartheta}_{1}\right)$ $\sim\mathcal{U},$ (128)
so under Conditions 1 - 3,
$\lim_{T\rightarrow\infty}\lim_{M\rightarrow T!}\hat{Q}_{M}=\lim_{T\rightarrow\infty}\lim_{M\rightarrow T!}\hat{\mathcal{H}}_{M}\left(\hat{\vartheta}_{1}\right)\xrightarrow{d}\mathcal{U}.$ (129)
The proof of Theorem 4 follows from the result of Theorem 1.
Proof of Theorem 4.
Under the alternative, we have $\vartheta_{1}<\vartheta_{i}$ for all possible permutations $i=1,\ldots,T!$. We observe $M$ permutations of the total possible. Define $\delta_{M}$ as the minimum difference between the underlying parameter $\vartheta_{1}$ and another permutation in $m=1,\ldots,M$, and $\eta_{M}$ as the maximum estimation error over all permutations $m=1,\ldots,M$.
$\delta_{M}$ $=\min_{1<m\leq M}\left|\vartheta_{m}-\vartheta_{1}\right|\geq\delta\quad\text{where }\delta=\lim_{M\rightarrow T!}\delta_{M}$ (130)
$\eta_{M}$ $=\max_{1\leq m\leq M}\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|\leq\eta\quad\text{where }\eta=\lim_{M\rightarrow T!}\eta_{M}$ (131)
If $2\eta_{M}<\delta$, then the quantile estimate $\hat{Q}_{M}$ will return the true value in the limit. Writing the minimum using the definition in Equation 9,
$\min_{2\eta_{M}<\delta}\hat{Q}_{M}$ $=\min_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\hat{\vartheta}_{m}\leq\hat{\vartheta}_{1}\right\}$ (132)
$=\min_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\hat{\vartheta}_{m}-\hat{\vartheta}_{1}\leq 0\right\}$ (133)
$=\min_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{(\hat{\vartheta}_{m}-\vartheta_{m})+(\vartheta_{1}-\hat{\vartheta}_{1})+(\vartheta_{m}-\vartheta_{1})\leq 0\right\}.$ (134)
The quantity in Equation 134 will reach its minimum when the estimation errors are maximized, leading to a larger value inside the indicator and fewer pairs that meet the criteria.
$\min_{2\eta_{M}<\delta}\hat{Q}_{M}$ $\geq\min_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{2\eta_{M}+(\vartheta_{m}-\vartheta_{1})\leq 0\right\}$ (135)
$\geq\min_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{2\eta_{M}+\vartheta_{m}\leq\vartheta_{1}\right\}$ (136)
For the maximum,
$\max_{2\eta_{M}<\delta}\hat{Q}_{M}$ $=\max_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{(\hat{\vartheta}_{m}-\vartheta_{m})+(\vartheta_{1}-\hat{\vartheta}_{1})+(\vartheta_{m}-\vartheta_{1})\leq 0\right\}$ (137)
$\leq\max_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{-2\eta_{M}+(\vartheta_{m}-\vartheta_{1})\leq 0\right\}$ (138)
$\leq\max_{2\eta_{M}<\delta}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\vartheta_{m}\leq\vartheta_{1}+2\eta_{M}\right\}.$ (139)
To show consistent behavior with the original ordering, if $2\eta_{M}<\delta$ and $\vartheta_{m}\leq\vartheta_{1}$,
$2\eta_{M}+\vartheta_{m}\leq 2\eta_{M}+\vartheta_{1}-\delta$ $\leq\vartheta_{1}$ (140)
$\quad\text{and}\quad\vartheta_{m}\leq\vartheta_{1}-\delta+2\eta_{M}$ $\leq\vartheta_{1}+2\eta_{M},$ (141)
or if $2\eta_{M}<\delta$ and $\vartheta_{m}>\vartheta_{1}$,
$2\eta_{M}+\vartheta_{m}>2\eta_{M}+\vartheta_{1}+\delta$ $>\vartheta_{1}$ (142)
$\quad\text{and}\quad\vartheta_{m}>\vartheta_{1}+\delta$ $>\vartheta_{1}+2\eta_{M}.$ (143)
Thus, the ordering is preserved.
$\lim_{M\rightarrow T!}\min_{2\eta_{M}<\delta}\hat{Q}_{M}$ $\geq\lim_{M\rightarrow T!}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\vartheta_{m}\leq\vartheta_{1}\right\}\geq Q$ (144)
$\lim_{M\rightarrow T!}\max_{2\eta_{M}<\delta}\hat{Q}_{M}$ $\leq\lim_{M\rightarrow T!}\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\left\{\vartheta_{m}\leq\vartheta_{1}\right\}\leq Q$ (145)
$\implies\lim_{M\rightarrow T!}\hat{Q}_{M}=Q\quad\text{when $2\eta_{M}<\delta$}$ (146)
We now need to show the probability of $2\eta_{M}<\delta$ goes to one in the limit. Pick any $\delta>0$.
$\mathbb{P}\left(\eta_{M}<\frac{\delta}{2}\right)$ $=\mathbb{P}\left(\max_{1\leq m\leq M}\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ (147)
From the result of Theorem 1, when $\tau_{\omega}^{2}=0$,
$\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ $=1.$ (148)
Similarly,
$\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ $=1.$ (149)
Returning to the maximum,
$\mathbb{P}\left(\max_{1\leq m\leq M}\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ $=\mathbb{P}\left(\left|\hat{\vartheta}_{1}-\vartheta_{1}\right|<\frac{\delta}{2},\ldots,\left|\hat{\vartheta}_{M}-\vartheta_{M}\right|<\frac{\delta}{2}\right)$ (150)
$\geq 1-\sum_{m=1}^{M}\mathbb{P}\left(\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|\geq\frac{\delta}{2}\right),$ (151)
and taking the limit under the two scenarios,
$\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\max_{1\leq m\leq M}\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ $\geq 1-\sum_{m=1}^{M}\lim_{\mathscr{R}\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|\geq\frac{\delta}{2}\right)=1,$ (152)
when $\tau_{\omega}^{2}=0$, and
$\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\max_{1\leq m\leq M}\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|<\frac{\delta}{2}\right)$ $\geq 1-\sum_{m=1}^{M}\lim_{\varphi\rightarrow\infty}\lim_{T\rightarrow\infty}\mathbb{P}\left(\left|\hat{\vartheta}_{m}-\vartheta_{m}\right|\geq\frac{\delta}{2}\right)=1.$ (153)
This implies $\hat{Q}_{M}\xrightarrow{p}Q$ in both cases presented in the statement of Theorem 4. ∎
9.2.1 Finite Sample Properties
For a finite sample, we can derive the distribution of the estimated quantity $\hat{\vartheta}_{m}$ as a sum of linear combinations of chi-square random variables.
Theorem 5.
With a finite sample containing $T$ observations, $\mathscr{R}$ random initializations, and $\varphi$ realizations in the set $\Omega_{\text{obs}}$, the estimate for the variation parameter $\hat{\vartheta}_{m}$ defined in Equation 6 follows a generalized chi-square distribution of the following form
$\hat{\vartheta}_{m}$ $\sim\frac{1}{\varphi\mathscr{R}K}\sum_{\omega=1}^{\varphi}\sum_{r=1}^{\mathscr{R}}\sum_{k=1}^{K}\frac{1}{T_{k}}\sum_{i=1}^{T_{k}d}\bigg[\left(\phi_{m,\omega,r,k,i^{\prime}i^{\prime}}+1\right)s_{m,\omega,r,k,i^{*}i^{*}}X_{m,\omega,r,k,i}$
$\hskip 36.135pt+\sum_{j=1}^{d\lfloor(i-1)/d\rfloor}\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}s_{m,\omega,r,k,i^{*}j^{*}}\left(Y_{m,\omega,r,k,ij}-Z_{m,\omega,r,k,ij}\right)$
$\hskip 36.135pt+\sum_{j=d\lfloor(i-1)/d\rfloor+1}^{i-1}\left(\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}+1\right)s_{m,\omega,r,k,i^{*}j^{*}}\left(Y_{m,\omega,r,j,ij}-Z_{m,\omega,r,k,ij}\right)\bigg],$ (154)
where $X_{m,\omega,r,k,i},Y_{m,\omega,r,k,ij},Z_{m,\omega,r,k,ij}\sim\chi_{1}^{2}$ for all $\omega\in\Omega_{\text{obs}}$, $r=1,\ldots,\mathscr{R}$, $k=1,\ldots,K$, $i=1,\ldots,T_{k}d$ and $j<i$, $\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}=(\phi_{m,\omega,r,k,ij})$, $\mathbf{S}_{m,\omega,r}=(s_{m,\omega,r,ij})$, $i^{\prime}=\lceil i/d\rceil$, $j^{\prime}=\lceil j/d\rceil$, $i^{*}=i\bmod d$, and $j^{*}=j\bmod d$.
The proof of Theorem 5 follows a step by step derivation of the distribution from the known starting point of Condition 3.
Proof of Theorem 5.
We begin with the distribution of the model residuals that is known from Equation 11 and Condition 3, $\mathbf{U}_{m,\omega,r}\sim\mathcal{MN}_{T\times d}\left(\mathbf{0},\mathbf{I}_{T},\mathbf{S}_{m,\omega,r}\right)$. The predicted residuals can be expressed as a function of the model residuals, like in Equation 5, and their distribution follows from that of $\mathbf{U}_{m,\omega,r}$ as in Equation 35. Like in the proof of Lemma 1, we define
$\bm{\Phi}_{m,\omega,r,k}=\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime},$ (155)
and further define
$\mathbf{V}_{m,\omega,r,k}\sim\mathcal{MN}_{T_{k}\times d}(\mathbf{0},\mathbf{I}_{T_{k}},\mathbf{I}_{d})$ (156)
such that we can write $\mathbf{R}_{m,\omega,r,k}=\left(\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right)^{1/2}\mathbf{V}_{m,\omega,r,k}\mathbf{S}_{m,\omega,r}^{1/2}$. The distribution has an equivalent vectorized form, with covariance matrix the Kronecker product between the row-wise and column-wise variation.
$\mathbf{r}_{m,\omega,r,k}=\text{vec}\left(\mathbf{R}_{m,\omega,r,k}\right)$ $\sim\mathcal{N}_{T_{K}d}\left(\mathbf{0},\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\otimes\mathbf{S}_{m,\omega,r}\right)$ (157)
For simplicity, define $\bm{\Delta}_{m,\omega,r,k}=\left[\bm{\Phi}_{m,\omega,r,k}+\mathbf{I}_{T_{k}}\right]\otimes\mathbf{S}_{m,\omega,r}$ and write the equivalent form $\mathbf{r}_{m,\omega,r,k}=\bm{\Delta}_{m,\omega,r,k}^{1/2}\mathbf{v}_{m,\omega,r,k}$, where $\mathbf{v}_{m,\omega,r,k}\sim\mathcal{N}T_{k}d\left(\mathbf{0},\mathbf{I}\right)$. The covaraince matrix $\bm{\Delta}_{m,\omega,r,k}=\left(\delta_{m,\omega,r,k,ij}\right)$ is symmetric and each element in the vector $\mathbf{v}_{m,\omega,r,k}=\left(v_{m,\omega,r,k,i}\right)\sim\mathcal{N}(0,1)$.
$\mathbf{r}_{m,\omega,r,k}^{\prime}\mathbf{r}_{m,\omega,r,k}$ $=\mathbf{v}_{m,\omega,r,k}^{\prime}\bm{\Delta}_{m,\omega,r,k}\mathbf{v}_{m,\omega,r,k}$ (158)
$=\sum_{i=1}^{T_{k}d}\bigg(\delta_{m,\omega,r,k,ii}v_{m,\omega,r,k,i}^{2}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{T_{k}d}\delta_{m,\omega,r,k,ij}v_{m,\omega,r,k,i}v_{m,\omega,r,k,j}\bigg)$ (159)
$=\sum_{i=1}^{T_{k}d}\bigg(\delta_{m,\omega,r,k,ii}v_{m,\omega,r,k,i}^{2}+2\sum_{j=1}^{i-1}\delta_{m,\omega,r,k,ij}v_{m,\omega,r,k,i}v_{m,\omega,r,k,j}\bigg)$ (160)
$=\sum_{i=1}^{T_{k}d}\bigg(\delta_{m,\omega,r,k,ii}v_{m,\omega,r,k,i}^{2}+\frac{1}{2}\sum_{j=1}^{i-1}\delta_{m,\omega,r,k,ij}\left(v_{m,\omega,r,k,i}+v_{m,\omega,r,k,j}\right)^{2}$
$\hskip 72.26999pt-\frac{1}{2}\sum_{j=1}^{i-1}\delta_{m,\omega,r,k,ij}\left(v_{m,\omega,r,k,i}-v_{m,\omega,r,k,j}\right)^{2}\bigg)$ (161)
We note that
$v_{m,\omega,r,k,i}^{2}$ $\sim\chi_{1}^{2}$ (162)
$\frac{1}{2}(v_{m,\omega,r,k,i}+v_{m,\omega,r,k,j})^{2}$ $\sim\chi^{2}_{1}$ (163)
$\frac{1}{2}(v_{m,\omega,r,k,i}-v_{m,\omega,r,k,j})^{2}$ $\sim\chi^{2}_{1},$ (164)
and $\delta_{m,\omega,r,k,ij}$ depend on the matrices $\bm{\Phi}_{m,\omega,r,k}=(\phi_{m,\omega,r,k,ij})$ and $\mathbf{S}_{m,\omega,r}=(s_{m,\omega,r,ij})$. Let $i^{\prime}=\lceil i/d\rceil$, $j^{\prime}=\lceil j/d\rceil$, $i^{*}=i\bmod d$, and $j^{*}=j\bmod d$. We define the following random variables for all $\omega\in\Omega_{\text{obs}}$, $r=1,\ldots,\mathscr{R}$, $k=1,\ldots,K$, $i=1,\ldots,T_{k}d$ and $j<i$.
$X_{m,\omega,r,k,i},Y_{m,\omega,r,k,ij},Z_{m,\omega,r,k,ij}$ $\sim\chi_{1}^{2}$ (165)
We can write each $\delta_{m,\omega,r,k,ij}=\left(\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}+\mathbf{1}\left\{i^{\prime}=j^{\prime}\right\}\right)s_{m,\omega,r,k,i^{*}j^{*}}$ and input the definitions.
$\mathbf{r}_{m,\omega,r,k}^{\prime}\mathbf{r}_{m,\omega,r,k}$ $=\sum_{i=1}^{T_{k}d}\bigg[\left(\phi_{m,\omega,r,k,i^{\prime}i^{\prime}}+1\right)s_{m,\omega,r,k,i^{*}i^{*}}X_{m,\omega,r,k,i}$
$\hskip 36.135pt+\sum_{j=1}^{d\lfloor(i-1)/d\rfloor}\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}s_{m,\omega,r,k,i^{*}j^{*}}\left(Y_{m,\omega,r,k,ij}-Z_{m,\omega,r,k,ij}\right)$
$\hskip 36.135pt+\sum_{j=d\lfloor(i-1)/d\rfloor+1}^{i-1}\left(\phi_{m,\omega,r,k,i^{\prime}j^{\prime}}+1\right)s_{m,\omega,r,k,i^{*}j^{*}}\left(Y_{m,\omega,r,j,ij}-Z_{m,\omega,r,k,ij}\right)\bigg]$ (166)
To obtain the distribution for the variation parameter, we aggregate this generalized chi-square distribution over the test sets $k=1,\ldots,K$, initializations $r=1,\ldots,\mathscr{R}$, and potential realizations $\omega\in\Omega_{\text{obs}}$. ∎